### v3

# Game-based Platforms for Artificial Intelligence Research

Abstract
Games have been the perfect test-beds for artificial intelligence research for the characteristics that widely exist in real-world scenarios. Learning and optimisation, decision making in dynamic and uncertain environments, game theory, planning and scheduling, design and education are common research areas shared between games and real-world problems. Numerous open-source games or game-based environments have been implemented for studying artificial intelligence. In addition to single- or multi-player, collaborative or adversarial games, there has also been growing interest in implementing platforms for creative design in recent years. Those platforms provide ideal benchmarks for exploring and comparing artificial intelligence ideas and techniques. This paper reviews the game-based platforms for artificial intelligence research, provides guidance on matching particular types of artificial intelligence with suitable games for testing and matching particular needs in games with suitable artificial intelligence techniques, discusses the research trend induced by the evolution of those platforms, and gives an outlook.
Impact Statement:
Games have been playing an essential role in the development of AI techniques and education by serving as valuable test-beds. This paper provides a comprehensive review of recent games and game-based platforms for both game playing and design. The platforms discussed in this paper are comprehensively categorised based on factors such as game type, aim, number of agents, observability, programming languages, and the presence of competitions. Furthermore, the paper investigates the challenges posed by games, existing issues with game-based platforms, and the efforts in leveraging natural languages for game research. By offering a detailed roadmap and research insights, this paper serves as a valuable resource for students and researchers, facilitating their exploration of this field. Besides, the paper presents the evolution of games along with the advancement of AI techniques, thereby igniting inspiration for future investigations. Readers from industry will also discover how AI can impact games and how games can impact AI research through this survey.
Index Terms:
Game AI, platform, reinforcement learning, procedural content generation, competition
I Introduction
Figure 1: Roadmap of games and platforms, created by authors using screenshots of the games and platforms and figures of papers cited in this work.
Games have played an essential role in the rapid development of artificial intelligence (AI). As a controllable, cost-effective, and realistic simulator [1], games provide a virtual world for AI to explore. A variety of real-life problems, such as real-time decision making, scheduling, design and education, can be reflected in such virtual worlds. Games have naturally become the perfect test-beds to verify and explore the ability, generality, robustness and safety of AI. Game-based platforms have provided an ideal playground for AI researchers to study, explore, evaluate and experiment with different ideas in a controllable and safe environment. Such platforms also enable us to study and answer what-if questions that are hard to do safely in the real world. They can even stimulate researchers to ask new research questions.
Various games have been used in game-based platforms, from single-player to multi-player, and from puzzles to real-time strategy games, with the emergence of diverse game genres and mechanisms. Such diversity produces a large number of diverse challenges for AI to solve. For instance, it has been demonstrated in various games (e.g., Chess [2], Atari [3], Poker [4], Go [5], StarCraft II [6], Dota 2 [7], and Texas hold’em [8]) that AI can perform at a level beyond that of humans. Besides playing games by AI, designing new games by AI is an emerging research area which helps exploit AI’s potential in creativity [9]. Games are rich in their content types, including visuals, audio, levels, rules, etc., making them challenging and diverse environments for innovative AI research [10, 11].
In terms of creative game design, search-based methods, machine learning (ML) and deep learning (DL) techniques have been used for procedural content generation (PCG) where both AI creativity and human-AI co-creativity are explored [12, 13, 14, 15, 16]. Emergent abilities of large language models (LLMs) also open up new opportunities and paradigms of playing and designing games [17, 18, 19, 20, 21, 22]. More recently, Bruce et al. from DeepMind introduce Genie, the generative interactive environment that is capable of learning from videos to generate playable, endless 2D platformer games from an image prompt [23]. AI-generated contents can help enhance the generality of AI itself [11]. AI techniques can help to automate the design tasks to accelerate the process, generate novel and diverse contents, and personalise contents for individual players.
This paper provides a bird’s-eye view of publicly available game-based platforms or games for AI research, aiming to provide a brief but comprehensive overview for researchers and students that want to either use AI for games or use games for AI. Fig. 1 presents the roadmap of games and platforms. Links to the source code of the reviewed games and platforms, corresponding research papers, and competitions are provided on the GitHub page [^1] All the links to the reviewed games, platforms and user instructions are collected at https://github.com/SUSTechGameAI/GameAIPlatforms..
The paper is organised as follows. Section II explains the scope and methodology of this survey. Taxonomies of playing and designing tasks involving games for AI research are provided in Sections III and IV, respectively. The research trends and challenges induced by the emergence and evolution of those game-based platforms are discussed in Section V. Finally, Section VI concludes.
II Methodology of Survey
Our survey considers the interpretation of games as Liu et al. [15], “any games a human would conceivably play, including board games, card games, and any type of video games, such as arcade games, role-playing games, first-person shooters, puzzle games, and many others”. Simulators for autonomous driving or robotic control are not covered since they are specifically designed for simulating and studying robotic systems with accurate physics simulations.
To simplify the language, from now on, we use the term “game” and “platform” to refer to a game simulator and an environment with multiple games for AI research, respectively. Yannakakis and Togelius [10] have enumerated and discussed some game-based platforms for game playing and design till 2018. Shao et al. [24] have reviewed some video game platforms for researching deep reinforcement learning till 2019. The work of Swiechowsk [1], Giannakos et al. [25], Duan et al. [26] have also touched on this topic. However, to our best knowledge, no existing paper has provided a complete review or taxonomy on recent game-based platforms, while more than 56 new ones have been released since 2018.
The following survey methodology is applied. First, we have gone through the competitions organised at related conferences in AI or game AI, including NeurIPS/NIPS, IJCAI, AAAI, ICML, ECML, ACML, AIIDE, IEEE CIG, IEEE CoG, ACM FDG, ACG, GECCO, WCCI/CEC, between 2016 and February 2024, and collected corresponding platforms. The year 2016 is chosen as AlphaGo [27] has boosted the research interest in game AI. Then, we searched with Google Scholar and Web of Science using the following search terms, “game” AND (“artificial intelligence” OR “AI”) AND (“design" OR “generation”) AND (“platform” OR “interface” OR “application” OR “software”), for related papers published or available online between 2016 and February 2024, assuming that most of the earlier ones can be found in Yannakakis and Togelius [10], Shao et al. [24], and then investigated if the papers used any game-based platform for AI research. Finally, for all related papers that did not specify the link to the corresponding platform or code, we searched for the code on the authors’ personal pages, in GitHub, and using Google Search. Again, only publicly available ones are included in our survey. It’s worth mentioning that a platform or game may have been available online before the corresponding paper is published. Besides, although some commercial games are not open-source, they provide some APIs to access the games for research, e.g., Starcraft II [6]. Such games are also included.
III Playing Games by AI
Games considered in this survey can be briefly categorised into tabletop games and video games. Tabletop games, including puzzles, board games and card games, usually have few requirements for image rendering, use simple objects such as boards, pieces and cards, and do not rely on digital devices. States of such games can be easily represented by structured data. Video games encompass a wide range of genres such as role-playing games, shooter games, racing games, puzzle games, and more. Unlike tabletop games, video games contain visual and audio contents based on digital devices. This poses a challenge of extracting high-dimensional but uninformative states and action features. Moreover, credit assignment in games may be challenging due to the game length, unobservable behaviours of players, combinational actions and sparse rewards. Tab. I, Tab. II and Fig. 3 categorise the games according to the game genre and the simulated physical space with which avatar(s) interact. Platforms equipped with multiple games are also introduced.
TABLE I: 36 games for playing by AI. The column with header C indicates whether an AI competition is associated. Geometry Friends, available before 2016, is included as it’s used in recent research. “1”, “2” and “>2” denote the number of agents, representing single-, two- and multi-agent. , and denote whether the platform is deterministic, stochastic or both of them. $*{}^{*}$ highlights commercial game with publicly available API. “A”, “C”, and “M” denote adversarial, cooperative, competing as mixed scenarios.
Type Platform Aim # Agents Observability Language Competition Reference
Planning Learning $1$ $2$ $>2$ Full Partial
Tabletop Games (Section III-A) Board ColorShapeLinks A C# [28]
Blood bowl A Python [29]
Mjai A JavaScript [30]
Mjx A Python [31]
Diplomacy M Python [32]
Card ben M Python [33]
Sabberstone A C# [34]
VGC A Python [35]
Tales of Tributes A C# [36]
Hanabi A M Python [37]
Social The Chef’s Hat A Python [38]
AI wolf M Java/Python [39]
Text Light M M Python [40]
Video Games (Section III-B) Grid- based 2D SimCity Python [41]
Nethack Python [42]
Snake AI A Java [43]
microRTS A Java [44]
Tribe A A Java [45]
Overcooked C Python [46]
Cuisine World C C++ [47]
Pommerman M Java/Python [48]
Neural MMO M Python [49]
Pixel- based 2D Mario AI Java [50]
MsPacMan-vs-Ghosts Java [51]
AIBIRDS Java/Python [52]
XFC Python [53]
Geometry Friends C C# [54]
AIGD2 A Java [55]
FightingICE A Python [56]
Honor of King $*{}^{*}$ A Python [57]
Dota 2 $*{}^{*}$ M Python/ Lua [7]
StarCraft II $*{}^{*}$ A M C++/Python [6]
3D ViZDoom Python [58]
MineRL Python [59]
MineDojo Python [60]
Hide and Seek M Python [61]
 Single-player deterministic game  Single-player stochastic game Configurable    $*{}^{*}$ Commercial game with API
Games with #Agent2absent2$\geq$ 2≥ 2: “A” for adversarial “C” for cooperative    “M” for mixed
TABLE II: 20 platforms with multiple games for playing by AI (cf. caption of Tab. I for reading instructions). Whether a game is adversarial or collaborative is not indicated as each platform contains a number of different games. GGP, available before 2016, is included as it’s widely used in recent research. “1”, “2” and “>2” denote the number of agents, representing single-, two- and multi-agent. Action-adventure game (AAG) involves video games such as puzzle and combat games.
Platform Aim Game genre #Agents Observability Language Competition Reference
Planning Learning Board Card AAG RTS Sport $1$ $2$ $>2$ Full Partial
GGP Java [62]
Botzone Java/Python [63]
OpenSpiel C++/Python [64]
Ludii Java/Python [65]
TAG Java [66]
Pgx Python [67]
Gym Python [68]
PettingZoo Python [69]
Minigrid Java/Python [70]
GVGAI Java/Python [71]
GVGAI gym Java/Python [72]
Stratega C++/Python [73]
Griddly Python [74]
Melting Pot Python/ Lua [75]
Retro Python [76]
Procgen Python [77]
Olympics-Integrated Python [78]
DeepMind Lab Python [79]
ML-agents Python [80]
MiniWorld Python [81]
III-A Tabletop Games
Board, card, puzzle games and other tabletop games raise several challenges to AI, such as partial observability and multi-agent cooperation or competition.
In ColorShapeLinks [28], players win by connecting pieces with the same colour or shape. The platform of ColorShapeLinks not only provide a visual front-end via the Unity game engine, but also a pure text-based one. ColorShapeLinks has been used for education and held as an AI competition, thus it can be easily accessed by individuals with different backgrounds. However, ColorShapeLinks is relatively simple for research purposes since good solutions can be quickly found. Blood bowl [29] simulates a fully-observable board game that two players control their own teams and are rewarded with touchdowns. At each turn, the player moves their avatars to block the team of the opposing player or pass the ball. The environment is compatible with Gym’s API for RL research [29]. In Blood bowl, the observation space consists of both spatial and non-spatial information and the action space differs across states, which leads to a new challenge in representation. Mjai [30] and Mjx [31] are partial-observable AI environments of riichi Mahjong. Thirteen tiles on the own hand and discarded cards are available for the player. If a special permutation of cards is formed, the player wins the game. Mjx is 100 times faster than Mjai in running time and is compatible with Gym’s API [31]. Diplomacy is a seven-person board game, considering both cooperation and competition [32]. In this game, AI agents are required to understand human players’ imitations and behaviours.
Hearthstone: Heroes Of Warcraft is a popular commercial turn-based collectable card game (CCG). An open-source implementation version called Sabberstone [34] is released for playing and deck construction. Pokémon Video Game Championships (VGC) is a turn-based strategy game [35]. Players need to collect Pokémons with different abilities to build a team. During the battles, player agents decide the order of Pokémons and specify the skills to attack the opponents according to previous information. The work of Budzki et al. [36] introduces a deck-building card game, Tales of Tribute. Players build the deck with the same initial cards during the game through battles and trades, meeting some dynamics in a long-term strategy. Tales of Tribute also provides a graphic interface and supports implementing AI agents with C#. The three games are all CCG, where players build decks and fight against opponents, but they have different research purposes. Sabberstone and Tales of Tribute both focus on the decision agents, i.e., how to play cards. However, Tales of Tribute also considers building decks, while Sabberstone doesn’t. Meanwhile, VGC involves meta-game balance that adjusts attributes of cards during the battles. Different from some traditional card games, players of Hanabi [37] can only see the cards of other players. During the gameplay, the player can tell others some hints about their cards. As a fully cooperative game, players need to form fireworks that consist of the same colour cards in ascending order to win. Besides Poker [4] and Texas hold’em [8], Bridge is another challenging board game for AI because of its partial observability, high branching factor, cooperation and competition involved in the game [82]. 23 editions of the World Computer-Bridge Championship have been organised with the latest one in 2019222World Computer-Bridge Championship: https://bridgebotchampionship.com/. The game engine of the championship was not found by authors. However, there is a recently released game engine called ben [33] for AI research in Bridge.
In most games, agents directly control the avatars to move or play cards, while in some others, agents take non-direct actions such as languages. The social card game, The Chef’s Hat [38] is also implemented based on Gym [68] and used for RL research. The winning goal of the game is to discard all hand cards [38]. AI wolf [39] is a multi-agent communication game that aims to find the wolf by talking with players. The coordination of multiple agents in partial observation only by talking is considered in the platform. Light is a text-based adventure game [40], with which dialogue-based human-AI interaction is studied.
Tabletop games usually don’t require specific designs in state representation, while the action space in those games changes during an episode. For example, in Go, the state is easily constructed by a matrix, where three symbols or numbers denote white, black or empty. However, the empty space for placing stones changes in a decreasing trend, leading to a variable action space. On the other hand, tabletop games often involve multiple players with complex and even dynamic relationship (collaboration vs. competition), pose new challenges and encourage studies in various research fields such as multi-agent systems [83], game theory [84] and psychology [85].
III-B Video Games
This section categorises the video games into grid-based 2D, pixel-based 2D and 3D according to the simulated physical space with which avatar(s) interact. Grid-based games are usually constructed by some preset tiles where players act with a set of discrete actions. Unlike grid-based games, pixel-based games feature continuous state space with either continuous or discrete action space. The difference in the state representation results in different feature extraction between grid-based and pixel-based games. The state in grid-based games can be easily converted to a low-dimensional matrix since the number of tiles is usually countable. On the other hand, feature extraction for pixel-based games is usually applied to states represented by high-dimensional images. 3D games extend 2D games with an additional height dimension, which not only enriches the visual and interactive depth of games but also introduces significant challenges in representing the state effectively. Besides, the action spaces of 3D games are usually more complicated or larger than those of 2D games.
The transition from grid-based 2D to pixel-based 2D and 3D games marks a significant evolution in video games during the development of technology. This progress poses new challenges and opportunities for AI research. As games become more sophisticated and closer to the real-world, AI techniques are evolving to meet these challenges simultaneously. Furthermore, the difference between games and 3D virtual environments for robotics and autonomous driving, such as MuJoCo [86] and CARLA [87], is narrowing.
III-B1 Grid-based 2D Games
Classic single-player and multi-player competitive games have been implemented using grid-based maps. SimCity [41] is an open-ended city-building 2D video game with an open-source version provided for reinforcement learning research. Nethack [42] is a single-agent roguelike game with procedurally generated levels. Agents meet the different stages at each episode, which drives AI’s abilities for long-term exploration and generalisation. Compared with existing single-player games, Nethack provides a faster simulator with massive tasks and game dynamics. Multi-agent games involve multiple independent decision agents, oriented towards cooperative and competitive behaviours. Snake AI [43] allows two snakes to fight for an apple. Since the apple is placed randomly in maps, the game can sometimes be unbalanced.
Modern strategy games incorporate digital elements to achieve complex and diverse mechanisms. microRTS [44] is a lightweight, grid-based real-time strategy (RTS) game framework that has fewer rendering sprites and focuses on the game’s mechanics, in which units are represented by different colours and shapes. Compared with StarCraft II, microRTS is easier to manipulate, so that researchers can focus more on the design of AI ideas and algorithms. As an implementation of The Battle of Polytopi, Tribe [45] involves multiple multi-agent scenarios as a turn-based strategy game framework, in which a forward model is supported for decision. Partial observability mode is supported by the framework, where the unexplored area is covered by the fog of war.
The work of Carroll et al. [46] provides a simple implementation of the commercial game Overcooked! for human-AI coordination, in which agents and players work together to cook meals. Agents and players need to move discretely and avoid colliding with each other, which introduces a new challenge of making motion coordination strategies. However, this implementation [46] encodes the game map as a 2D array, which can hardly express the complete information of Overcooked!. For instance, the representation of the game level can not distinguish whether an onion is on a table or the floor. Besides, the ingredients are fixed in the implementation. It’s worth mentioning that the work of [47] also introduces a game engine similar to Overcooked!, called CuisineWorld, to evaluate multi-agent cooperation performance. The platform has been tested by LLM agents [47]. Though CuisineWorld is a 3D game engine, to our best knowledge, 2D game observations are used in [47]. Instead of traditional representation based on visual information, CuisineWorld inverts the game state with text-based representation and the action space consists of several instructions like moving and putting. Compared to the implementation by [46], CuisineWorld is more suitable for research on LLMs. Pommerman [48] is a multi-player competitive game for at least four players, who start with a limited number of bombs. Players move in a grid-based map and plant the bombs to destroy their opponents. Pommerman allows multiple players to cooperate and compete as a team on the battlefield. Neural Massive Multiplayer Online (Neural MMO) [49] considers a large multi-agent scenario where at most 1024 agents learn to cooperate and compete to survive in the grid-based world at the same time.
Grid-based games usually have simplified state and action space, as well as clear game mechanisms, which allow to train and test AI agents at a faster speed. However, this simplicity limits its usage in AI approaches considering continuous decision space and complex strategies.
III-B2 Pixel-based 2D Games
Pixel-based games support more complex game mechanisms with continuous state space, while new challenges are introduced.
Mario AI framework [50] is based on the game Super Mario Bros., where players control Mario to run and jump through some barriers in a partially observable environment to arrive at the endpoint. However, the provided levels are simple and a limited number of levels are available, planning algorithms such as $*{}^{*}$ can quickly solve it without leveraging sophisticated AI approaches. MsPacMan-vs-Ghosts [51] is the implementation of Ms. Pac-Man arcade game, in which the player controls Ms. Pac-Man to collect pills while avoiding being hunted by four ghosts. Both full and partial observability are supported in MsPacMan-vs-Ghosts. AIBIRDS [52] is the platform with the popular puzzle game Angry Birds. Agents are required to shoot pigs to destroy the building made of blocks of different types and shapes. XFC [53] simulates the game Asteroid Smashers, in which AI controls a spacecraft to avoid asteroids and destroy pesky rocks. XFC designs a specific agent architecture, which is suitable for research on explainable AI and fuzzy AI.
Geometry Friends [54] provides a cooperative scenario that two agents aim to solve puzzles, restricted by real physics conditions such as friction and gravity. A simple implementation of Planet Wars, named AIGD2 [55], brings a new challenge with extra-long skill-depth to RTS games. FightingICE [56] considers a 2D video fighting game that AI agents fight against opponents according to sound. “Blind" agents face challenges such as partial observability and insufficient training.
Honor of Kings Arena, Dota 2 and StarCraft II [6] are three popular commercial games with publicly available APIs that can be used for AI research. Honor of Kings Arena [57] provides an interface to interact with Honor of Kings 1v1, in which agents learn to fight against diverse opponents. An API of Dota 2 written with Lua language is available for investigating multi-agent RL [7]. Fig. 2 presents the heroes controlled by OpenAI Five [7]. An interface connecting to StarCraft II [6] is also released for research in multi-agent RL and map design. Source codes of commercial games are often inaccessible, which means AI researchers can hardly edit commercial games to meet some particular requirements for research, such as creating an abstract, faster forward model of the game, even though their APIs are sometimes available.
Figure 2: Screenshot of OpenAI Five [7].
Figure 3: Games and platforms for playing, categorised as board/card, grid-based 2D, pixel-based 2D, 3D, social and text game types.
III-B3 3D Video Games
3D video games provide a more realistic simulation of the real world. Decision agents need to take actions based on the 3D vision instead of a flat plane in 2D games, which requires effectively processing high-dimensional visual information to make effective decisions.
ViZDoom [58] is a highly customisable doom-based environment with partial observability. Raw visual screens from the first-person perspective are received by the agents. ViZDoom is lightweight and fast. Its integrated editor allows users to manually create diverse task suits. ViZDoom is suitable as the test-bed for solving visual RL problems.
MineRL [59] is based on Minecraft, where a large scale human dataset is provided. The work of Baker et al. [61] presents a specific platform for Hide and Seek, which can be used for training and understanding multi-agent cooperative and competitive AI. MineDojo [60] provides a Gym-like environment based on Minecraft. Agents are supported to directly access the game for multiple tasks including material mining, crafting and building. Besides, a massive multimodal dataset, captured from video websites, gameplays, wiki pages and Reddit posts is provided.
Figure 4: Summary of games and platforms used for studing AI for game design, created by authors using screenshots of the games and platforms, as well as figures from their corresponding papers, reviewed in Section IV.
III-C Platforms
Classic and modern board, card and puzzle games are often integrated into a single platform as the AI test-bed. For example, General Game Playing (GGP[62] is designed to evaluate AI’s ability in playing different kinds of tabletop games, where rules and states of games are defined with game description language, and has been widely used in AI research and education. Hundreds of single- and multi-player games, such as Go, Tic-Tac-Toe, reversi and minesweeper, are available in GGP. Botzone [63] focuses more on large multi-player games and is famous for its Mahjong AI Competition. Botzone also support 2- and 6-person Texas hold’em. Besides some popular classic games, OpenSpiel [64] provides Poker and Bridge, as well as some classic algorithms and visual tools. Ludii [88] is another general game-playing system based on similar principles to Ludi [89] but with different mechanisms. It includes various types of games such as board games, boardless games, stacking games, single- and multi-player games. Ludii uses ludemes as game description grammar. The description language is easy to read and understand according to [88]. Different from GGP and Botzone, the Tabletop Games Framework (TAG) [66] is designed to study general AI in modern tabletop games and has implemented more than 20 games, including Uno, Love Letter and Virus!. An extra analysis tool is equipped in TAG for studying action space, branching factors, hidden information, and so on. Pgx [67] is a environment of board games for RL research. It is written in JAX and can do parallel simulation via GPU/TPU. Pgx can hold single-/multi-agent board games and Atari-like environments including Shogi, Go, Chess, and Backgammon. According to [67], Pgx can perform 10-100 times higher throughput compared to other Python libraries.
OpenAI Gym [68] and PettingZoo [69] offer a number of grid-based and pixel-based 2D games. OpenAI Gym [68] provides an API standard for RL research and contains massive games, such as Atari, as well as a wide range of third-party environments. However, Gym typically focuses on the single-agent setting. To conduct research in multi-agent RL, PettingZoo [69] extends the API standard of Gym and integrates some famous environments like multi-player Atari games, classic card/board games and multi-agent particle environment [90]. As providing the standard API, Gym and PettingZoo are now the most used and compatible platforms, which enhance the evaluation and reproduction in the RL community with a higher speed.
Minigrid [70] is a single-agent platform compatible with OpenAI Gym. The included games are simple in rendering content but can implement complex mechanisms. General Video Games AI (GVGAI) [71] platform and GVGAI gym [72] offer a larger number of grid-based single- and two-player games, and has been used in its single-player planing, two-player planing, and single-player learning competitions. Users can create new games and levels in GVGAI in an intuitive way using Video Game Description Language (VGDL) [91]. Stratega [73] evaluates the generalisation of strategy agents with multiple games and levels with an $n$-player turn-based setting. Profiling and logging tools are integrated for analysis. Usually, grid-based platforms are highly scalable in terms of game mechanisms and levels, so researchers can easily design specific tasks to address their research problems with the help of grid-based game platforms. Griddly [74] particularly provides grid-based real-time strategy games with various numbers of players, with a faster running speed and lower memory than Minigrid and GVGAI. Melting Pot [75] is a platform with more than 50 diverse games for evaluating the generalisation of multi-agent RL.
The Retro Learning Environment [76] extends Gym, by incorporating popular video games from Super Nintendo Entertainment and Sega Genesis. Procgen [77] evaluates the efficiency and generalisation of RL with 16 procedurally generated Gym environments, the initial state of which can vary for each episode. Olympics-Integrated [78] poses the challenge of designing one single agent to participate in various two-player Olympics games. Since the integrated games in Retro, Procgen and Olympics-Integrated are pixel-based, challenges like feature extraction are introduced for AI agents. Besides, these pixel-based game platforms usually have lower scalability in adding new games and tasks than grid-based platforms such as GVGAI.
DeepMind Lab [79] acts as a 3D learning environment which is based on some open source software such as ioquake3. It provides a set of challenging 3D navigation and puzzle-solving tasks for agents to learn. The primary purpose of DeepMind Lab is to act as a working platform for results in AI, especially deep reinforcement learning. ML-agents [80] is an open-source platform that allows to create new games in a simpler and more flexible manner with the help of the Unity3D toolkit. MiniWorld [81] provides a more accessible way to construct 3D environments for reinforcement learning and robotics research. It is qualified to generate and simulate environments with rooms, doors, mazes and other objects. MiniWorld is simpler than DeepMind Lab and ViZDoom in the view of engineering, so beginners can get started quickly without mastering Unity3D.
III-D Guidance for Playing Games
According to the aim, number of agents, observability, language and competition, games and platforms reviewed in this work are categorised in Tabs. I and II. Those games and platforms are often developed for specific research purposes.
Card/board-related games and platforms, including Blood bowl [29], Sabberstone [34] and Botzone [63], are suitable for research on tree search, causality and competitive scenarios in multi-agent systems. GVGAI [71, 72, 92], Stratega [73] and Procgen [77] are developed for the purpose of evaluating the generalisation of AI agents. Both GVGAI and Stratega can be used for research on planning and learning, while Procgen specially focus on RL.
In the single-agent setting, Nethack [42] considers how an AI agent solves exploration, planning, and inventory management problems in a procedurally generated 2D environment. Both ViZDoom [58] and MineRL [59] require AI agents to explore the game with 3D observation and action space and support RL. ViZDoom has some clear goals and limited accessible area, while MineRL poses challenges that agents need to explore an open world for survival. For the multi-agent setting, Snake AI [43], microRTS [44], AIGD2 [55] and Tribe [45] provide test-beds for competitive scenarios. However, Snake AI and AIGD2 are limited to planning. Both [46] and [47] implement Overcooked! to study the cooperative behaviours of multiple AI agents. Nevertheless, the former considers how agents cooperate via multi-agent RL, while the latter applies the text-based representation in the game and aims to plan the high-level actions of agents such as moving by LLMs. Moreover, Neural MMO [49], Dota2 [7] and Hide and Seek [61] are suitable for learning-based AI in the mixed multi-agent setting with partial observability. Neural MMO also consider the large-scale case in terms of the number of agents, compared with the other two games.
Besides, the integrated tools differ between games and platforms. For example. in GVGAI, users can easily construct their games and levels via VGDL. Pommerman [48] supports highly self-defined maps. However, game levels in Procgen can only be generated randomly. It is recommended to get started from Gym [68] and PettingZoo [69] for single- and multi-agent RL research since they are the most common platforms. Some platforms such as DeepMind Lab [79] and ML-agent [80] usually are equipped with some visual tools and loggers to analyse the behaviour of agents, but the requirement to learn about the game engines like Unity is crucial.
From the two tables, Python is the most used programming language, which aligns with the trend of the famous Python libraries such as Torch and TensorFlow.
IV Designing Games by AI
TABLE III: 25 environments for designing games by AI. Although Mario-AI, HOG2, PhyDSL, StepMania, OSU! and MiniDungeons are available before 2016, they have been used for testing and evaluating the procedurally generated SMB levels in recent years.
Type Environment Language Competition Reference
Game Tuning (Section IV-A) Parameter FightingICE Java [93]
AIGD2 Java [55]
VGC Python [35]
Content Generation (Section IV-B) 2D level Mario AI Java [50]
HOG2 C++ [94]
Procedural Zelda Lua [95]
Talakat JavaScript [96]
Science Birds Java [97]
Morai-Maker-Engine C# [98]
PCGRL gym Python [99]
micro-rct Python [100]
Keiki Python [101]
3D level GDMC Python [102]
The Sims C++ [103]
Narrative TextWorld Python [104]
Evennia PostScript [105]
Rhythm chart StepMania C++ [106]
OSU! C# [107]
Card XMage Java [108]
Metastone Java [109]
Other (Section IV-C) Code PhyDSL Java [110]
Persona MiniDungeons Java [111]
Data set VGLC DOT [112]
Interactive Baba is y’all Python [113]
Lode Encoder JavaScript [114]
In recent years, game-based platforms have also been used to create new tasks for AI and to study AI creativity, in particular, the ability of creating new contents. Tab. III and Fig. 4 summarise the games and platforms used for game design. Those platforms provide diverse and representative test-beds for discovering the potentials in AI-involved game design.
IV-A Game Tuning
Game tuning is vital since it directly affects players’ experience. Unsuitable game parameter settings may lead to game unbalance or low quality, where players find it hard to play the game or feel bored. Generally, all the games introduced in Section III that involve configurable parameters can be tuned with search-based methods for different purposes, such as adjusting the game difficulty. For instance, FightingICE [56] has been used to study the dynamic difficulty and battle sound adjustment [115]. VGC [35] allows adjusting the attributes of Pokémons, called meta-game balance. Compared to other platforms, it specifically facilitates the investigation of enabling fair competition in gameplay. AIGD2 [55], an open-source game similar to Planet Wars, allows to explore the parameter space, which can significantly affect the gameplay via GVGAI agents [116].
IV-B Content Generation
Content generation is involved with applying AI techniques to create new game content. The generated content types include 2D level, 3D level, narrative, rhythm chart and card, following Liu et al. [15]. Different content types are represented in different data structures, raising various challenges to content generation research.
IV-B1 2D Map & Level
Game level is the most commonly generated content type. A 2D level is usually presented in a plane, where agents move in two dimensions. Mario-AI-Framework [50], a fully-Java copy of Super Mario Bros., provides a tile-based representation of game levels. Fig. 5 shows examples of Mario levels generated by [117], tested and visualised in Mario-AI-Framework [50]. Hierarchical Open Graph 2 (HOG2) [94] includes several grid-based puzzle games such as Snakebird [118] for exhaustive PCG [119]. Morai-Maker [98] is a Unity3D-based platform for Super Mario Bros., which provides three built-in machine learning-based 2D level generators. Co-creation between humans and AI is supported by the platform with a visible game level editor [98], which distinguishes Morai-Maker from other Super Mario Bros. platforms. Procedural Zelda [95] is a grid-based game platform for level, combat and puzzle generation. Both offline and online control are supported during the generation with data logging capabilities. PCGRL gym [99] implements RL environments for grid-based level generation with the Gym interface. Binary, Zelda and Sokoban problems are integrated into PCGRL gym. micro-rct [100] is a theme park management simulator, inspired by RollerCoaster Tycoon 2. In [100], researchers use a controllable PCGRL to generate game map. micro-rct is used to evaluate the controllability of PCGRL. Unlike other platforms which serve for search-based, supervised learning-based, or unsupervised learning-based content generation algorithms, PCGRL gym [100] provides a standardised benchmark for RL-based PCG, however, no accessible playing interface for human players is provided. Science Birds [97] provides a generative platform, which can load XML-based level files to play Angry Birds, a physics-based puzzle video game, which requires the level components’ physical feasibility and stability. Generating levels for physical-based puzzle games with AI is complex since more physical constraints, larger state space and action space should be considered. Angry Birds AI Level Generation Competition [52] uses Science Birds as its simulator.
Figure 5: Mario level segments generated by the work of [117] tested and visualised in Mario AI framework [50].
Recently, LLMs have obtained much attention [120, 121]. Chatgpt4PCG [122] competition aims at using ChatGPT to generate game levels of Angry Birds. Although no specific environment API is provided, an example prompt is introduced. Chatgpt4PCG enables level generation via natural language, which is more user-friendly for beginners. Meanwhile, the investigation into generating levels with ChatGPT can help researchers to better understand the ability and limitations of LLMs. Talakat [96] is a bullet hell game engine which defines a domain-specific language for bullet hell levels, supporting the grammar evolution of bullet hell game content. Keiki [101] is another bullet hell type game engine which provides a native Python interface describing bullet hell barrage. While Talakat represents barrages as scripts, Keiki provides a function to convert barrage scripts into vector-based representations, making it convenient for ML-based content generators.
IV-B2 3D Map & Level
3D maps and levels are closer to real-world scenarios but their generation is more complex. Due to the additional dimension, the data scale of 3D maps or levels is typically much larger than those of 2D ones, requiring higher computational efficiency. Moreover, 3D maps and levels remain more complex spatial structures, which lead to higher requirements in playability and diversity. The Settlement Generation Competition for Minecraft (GDMC) [123] invites participants to procedurally generate a city or village given a Minecraft map, and directly uses the original game engine for simulating and experiencing the output city or village. A code interface is provided to build Minecraft maps by setting blocks. Fig. 6 shows examples of entries to the 2023 edition of GDMC. The Sims [103] is a game-inspired platform for indoor environment generation. In The Sims, AI techniques can be used to generate furniture and room decorations.
Figure 6: Screenshots of cities generated by an entry [124] to the GDMC AI Settlement Generation Challenge in Minecraft [123] in 2023.
IV-B3 Narrative
AI techniques can be used to generate interactive fiction worlds with semantic coherence and fun. Distinguished from generating other types of content, narrative generation requires AI to understand temporal logic, context memory and natural language. WorldGeneration [105] integrates DL-based and rule-based models to generate graph-based world representations, and then generate narratives of the world. The used game platform is Evennia, a Python-based framework for texted-based multiplayer online games, also known as multi-user dungeon. Evennia has a convenient API and can be easily accessed by the web client. Textworld [104] is a text-based engine, which can generate and simulate text games. The aim of Textworld is to provide an accessible API for RL agents [125] to understand natural language and then make decisions.
IV-B4 Rhythm Chart
Rhythm game is a special 2D game that aims to hit charts at the right time. To optimise player experience, the generated charts need to be synchronous with the rhythm of the music. StepMania [106] is a free dance and rhythm game engine based on game Dance Dance Revolution that supports customised rhythm charts. StepMania has been researched in level generation i.e. chart generation [126], which is related to the music information retrieval field using long short-term memory (LSTM) and C-LSTM model. StepMania can be used in difficulty estimation [127] as well. A beatmap and rhythm chart can be generated automatically according to the music [128, 129] by a rhythm game platform OSU! [107].
IV-B5 Board & Card
Content generation for board and card games encompasses the rules, mechanism or deck building. Ludi [89] is a general game system, aiming at procedurally generating board game rules using evolutionary methods. CCG provides a flexible way to rebuild the game mechanism using AI techniques. CCGs are challenging in the field of AI due to their vast search space and dynamically changing game states resulting from a large number of diverse cards and their unique functionalities. The famous CCG, Hearthstone has some open-source simulators such as Metastone [109] and Sabberstone [34], which has been used for deck building [130], card generation [108] and game playing [109]. Magic: The Gathering is another commercial CCG, which has an open-source simulator called XMage for AI research including text design [131] and game playing [132].
IV-C Other Resources for AI-involved Game Design
PhyDSL provides an accessible environment to design physics-based 2D games via code generation [110]. A high-level description language is developed to describe complicated game dynamics in PhyDSL. MiniDungeons [111] is a roguelike dungeon platform for persona modelling, which can be used to build game evaluators. Video Game Level Corpus (VGLC) [112] is an open level set of video games, including Super Mario Bros., Kid Icarus, The legend of Zelda and Doom2, in tile-based, graph-based (using DOT language in Graphviz [133]) and vector-based representations. Although the codes of some platforms are not released, interactive interfaces are accessible to game designers, e.g., Baba is y’all [113] and Lode Encoder [114].
IV-D Guidance for Designing Games
We summarise the games and platforms for game design in Tab. III. A majority of game design platforms focus on content generation. Programming languages vary across the games and platforms, with half of them supporting Java. Considering the trend of an increasing portion of using Python in AI research, it is recommended to use those platforms with cross-language programming. For example, researchers can access Java programs in Python via Jpype [^3] https://github.com/jpype-project/jpype/.
Beginners are recommended to try 2D level generation first since it is the most studied with massive platforms. In particular, the Mario AI framework is a widely used and representative platform, which is proper to get started. Mario AI, Procedural Zelda and PCGRL gym can be used to generate 2D grid-based levels. Specifically, researchers interested in PCG via RL can use PCGRL gym, while the other two platforms are suitable for studying search-based and supervised learning-based level generating methods. To study the generation of structured game levels with physical property, Science Birds is an appropriate choice. While both Talakat and Keiki are designed for bullet hell games, the former serves for grammar evolution while the latter is convenient for ML research.
For other types of content, there are very limited numbers of platforms to use. Researchers can directly choose the corresponding platform according to the interested content type with the help of Tab. III. FightingICE, AIGD2 and VGC are platforms that can be used for studying game tuning and balancing, but with different focus on fighting games, RTS games and Pokémon, respectively.
V Discussion and Outlook
New platforms and games have been published over the last few years as ideal test-beds for AI research, including multi-agent RL, artificial general intelligence and creative AI. This section discusses the core challenges from games, our key findings and outlook.
V-A Challenges from Games
Although originally designed for human entertainment, games have proven to be ideal test-beds for validating the capabilities of AI techniques, providing a range of challenges for AI to overcome.
V-A1 Curse of High Dimensions
Dimensions of games are involved with the number of game states, representation complexity of state and action spaces. For example, Go has about $3^{361}$ possible positions, encoded as a 0-1 matrix. Modern games that rely on digital devices have a more sophisticated, high-dimensional representation, i.e., video frames. Traditional tabular methods such as Q-learning [134] fail to deal with the situation since the required space scales exponentially with the dimension. It is from Deep Q-learning [3] that the neural network appears on the scenes as the direct state estimator. In recent years, the development of games has resulted in fancy game screens with high-quality graphics, which can immerse players in more sophisticated virtual worlds such as Atari [68], Dota2 [7] and StarCraft II [6]. However, this visual sophistication makes it hard, particularly for AI agents to extract relevant features from those games [7]. Well-rendered sprites can obscure critical information that AI agents need to make decisions. For example, in a game where players control characters that move around in an environment, the sprites can prevent AI agents from identifying obstacles and enemies. Furthermore, video games often provide players with a large and complex set of actions, leading to a combinatorial explosion of possible actions [135].
V-A2 Poor Exploration
AI techniques such as RL usually need a large amount of interaction data to train policies. It is hard to directly access such a “dataset”. Instead, we need to collect the data, denoted as experience, during training, which leads to a tradeoff between exploration and exploitation. The tradeoff happens with a limited computational budget since we hardly go through all possible states and actions. Montezuma’s revenge epitomises the games where AI techniques suffer from the tradeoff [136], which brings a deep thought about the exploration strategy. Integrating the exploration signal into the reward function has been shown efficient to tackle the challenge, such as count-based exploration [137], intrinsic curiosity [138] and novelty search [139]. While the above reward shaping may lead to the optimal policy variance, population-based exploration [140] and noisy exploration [136, 141] that perturb the policy parameter space present their advantage for the diverse exploration.
V-A3 Partial Observability
While complete information is known in Atari [68] and Go [5] at a given state, partial observability is encountered in more complex video games such as ViZDoom [58]. There are two meanings associated with partial observability: agents cannot obtain perfect information about the current state, and agents cannot determine the tendencies of other agents. The latter usually refers to the multi-agent case. In the former aspect, agents only can partially observe the environment [142]. ViZDoom [58] only allows players to view the video frame from a first-person perspective. Although agents can handle situations in front of them, unexpected occurrences such as monsters approaching from behind should also be considered.
V-A4 Generalisation
Although AI techniques such as deep RL have achieved a competitive level beyond humans in games including Go [5] and Atari [68], their generalisation ability is questionable since they are usually only trained on a given game or level. The generalisation considers if the agent can still play well in some unseen games and levels without overfitting the training environment [143]. Many game-based environments are involved with the topic. GVGAI [71, 72] focus on the performance of agents when trained on some game levels but tested on unseen game levels. The unseen test game levels have different game maps or layouts, but their mechanism is the same [77]. Both planning and learning algorithms are supported by the GVGAI environment [71, 72]. Procgen [77] also provides diverse game scenarios, which test the AI agents by randomly changing the initial position of the avatars.
V-A5 Multi-agent Games
It is common to see the participation of multiple agents or players in games such as tabletop games [36, 28], multiplayer online battle arena games [7], and RTS games [6]. Three main multi-agent settings are involved: fully cooperative, fully competitive, and mixed. For example, in Dota2, five players are matched as a team to fight against another 5-player team. The five players cooperate and share their resources and experience, while also competing with the other team. This participation leads to an increase in dimension with the expanding number of agents. Additionally, non-stationarity and credit assignment problems arise when multiple independent agents exist since it is difficult to determine the unique contribution of each agent. Another interesting exploration is the coordination between AI agents and humans. While humans typically rely on natural language to communicate and cooperate, AI agents may not have the same capacity. The recent emergence of LLMs [144] such as ChatGPT [145] shows the potential for human-machine interaction. The work of [146] present how LLMs cooperate in 7-person Werewolf without tuning. LLM-based players collect experience in each game round and demonstrate strategic behaviours. It remains to be seen if this success can be replicated in games such as Hanabi [37] by both communicating and acting. Integrating LLMs with current AI technologies such as RL is a potential solution. In this combination, LLMs can understand the current game state and extract information, which enables an RL agent to make reasonable decisions.
V-B AI for Games
The traditional developing process of games is long and complex, which burdens humans in designing and programming game content. Nowadays, AI can assist humans to improve games by testing and generating new content.
V-B1 Game Testing by AI
Game quality assurance is crucial in the gaming industry for assessing game content [147]. However, employing human testers incurs significant financial and time costs. To solve this problem, automated game testing has been widely studied for decades. With AI technologies such as search-based methods [148] and RL [149, 150], bugs in games can be automatically detected and fixed. Yet, current AI testing methods mainly lack sufficient generalisation across different games. Typically, AI-based testing solutions require additional costs of being adapted to new games. For example, search-based methods need to search for new test cases and learning-based methods require retraining [149]. Developing a more generic testing framework with lower cost is invaluable. A promising direction may involve transfer learning [151], which enables the automated testing framework to adapt to new games, making high-quality assurance more accessible for developers across the game industry.
Moreover, games often require frequent updates to introduce new content and address gameplay issues. This necessity can be both a source of excitement and frustration for players. The critical challenge lies in testing updated content and maintaining game balance. Such studied [152, 153] have been conducted on multiple games such as VGC [35] and StarCraft II [6]. Advanced testing methodologies are capable of adapting to the rapidly changing landscape of game development, which enhances the player experience without compromising game quality or balance.
V-B2 Game Design by AI
In games, levels is usually the most fantastic part, which represents the effort and inspiration of game designers. However, creating levels requires a significant investment of time and human resources. AI has the potential of accelerating this process and can even serve as a substitute for human designers [12, 13, 14, 15, 16]. Besides, AI can be applied to generate music and images in games. For example, Ferreira et al [154]. proposed a DL-based music composition system to generate music for tabletop role-playing games. Hong et al. applied multi-discriminator generative adversarial networks to generate game sprites [155]. On the other hand, other 2D and 3D content generation methods [156], not limited to AI approaches, can also be tested and applied to games. Moreover, game stories and programming codes can be generated via AI technologies. For example, Ubisoft has recently proposed Ghostwriter to assist game designers to write game stories and dialogues [157]. Ghostwriter receives feedbacks from designers to construct conversations between NPCs, instead of replacing designers directly.
Given these advancements in AI technologies, an interesting question arises: is it possible for AI to generate a complete game? GameGAN proposed by NVIDIA [158] presents this possibility. After training with massive playing records in Pac-Man, GameGAN can simulate the game without relying on a game engine or coding. When the player pushes some buttons, GameGAN predicts the next game frame for the player. More recently, DeepMind has proposed Genie [23], a generative interactive environment directly trained by videos from Internet. Beyond GameGAN [158], Genie generates endless virtual worlds in a more controllable way by inputting prompts and photos, with which players can directly interact. However, Genie is limited to generating 2D games presently. Besides, the generated game frames can be hardly modified manually. It is believed that there will be more AIs capable of generating complete games, marking the dawn of a new era in game development.
V-C Games for AI
Beside the specific challenges such as high dimensions and partial observability, games can enhance AI by providing unique testing environments. Games not only expand AI capabilities, but also are regarded as preliminary experiments before applying AI to real-world scenarios.
V-C1 Open-world Games for Artificial General Intelligence
General (video) game playing has been extensively studied using multi-game platforms [62, 71, 72, 92]. However, the objectives or tasks in such games are often well-defined and known. Open-world games, e.g., MineRL [59], pose new challenges for AI to autonomously explore the Minecraft world without specific aims, including the requirement of accomplishing a series or set of tasks instead of one or two, with inefficient data sampling and indeterminate reward function [59]. The nature of open-world games provides an ideal platform to develop and evaluate artificial general intelligence. They can be considered as sandboxes, where AI agents are not only tested on some specific tasks such as navigating and building, but also are required to interact with the dynamic environment. The diverse tasks such as farming, hunting and crafting in Minecraft enable AI to live as humans. This case is pushing the boundaries of AI towards achieving artificial general intelligence, as they mimic the complexity and unexpected of real-world tasks and decision-making.
V-C2 Games as Reflection of Real-world Scenarios
Games can simulate real-world scenarios and enable us to make new attempts of AI safely at a low cost and reflect challenges in real-world scenarios. For example, urban planning can be found in games such as SimCity [41]. Building and sculpture design prototypes are found in GDMC [102]. AI wolf [39] benefits for understanding the behaviours of AI in cooperative and competitive settings. Logistics problems are challenging in the real world and it can be difficult to evaluate the real-time decision algorithms directly in real-world manufacturing systems due to some safety and property constraints. These problems can be easily modelled in games to validate the effectiveness and safety of algorithms. For example, Shapez  [^4] https://github.com/tobspr-games/shapez.io is a building game that aims to simulate manufacturing by placing and combining different belts and workstations. AI can test extensive possible solutions with the game at a low cost, which contributes to the design of a factory in the real world. It is notable there are also many games like Mini Metro, Mini Motorway and Cities: Skylines that reflect real-world scenarios. However, the codes of those commercial games are not released.
V-C3 Emerging Environments for AI Creativity
The environments collected in this paper indicate the research trend from intelligent decision-making to creative AI. Since 2016, 14 new environments have been published for game design by AI. However, most of them aim at designing tile-based levels. In addition, the collected environments are often not equipped with detailed descriptions or documentation as they were released as dependencies of a published algorithm or ML model. Although PCG approaches have been studied or applied to a wide range of games, the environments are rarely open-source (e.g., Cut the Rope [159]). Besides, how to design effective content-evaluation metrics to enable promising content generation is a non-trivial problem. More open-source environments with diverse genres and content types, user-friendly APIs and documents, are expected to advance the research in AI creativity and human-AI co-creativity.
V-D Under-explored Game Genres
Although there are many available games, they share a large number of similar mechanisms with only differing in sprite styles. Many platforms for AI research are based on classic games like Atari [68]. Although the simplified representation diminishes some engineering problems, this may not lead to a more delicate game mechanism thereby promoting AI research. Only a few modern commercial games such as StarCraft II [6], and Dota 2 [7], provide accessible APIs or lightweight simulators.
Extensive effort has to be put into solving engineering problems such as controlling avatars, which is a critical issue in conducting AI research on commercial games. A possible solution for this is capturing the screening as input. AI agent directly accesses the screen frames and plays the game via a virtual keyboard and mouse. As new games are developed, challenges arise for AI research. It is expected that more commercial games provide accessible APIs that AI can interact with. Similar concerns exist in game-based platforms as most platforms for content generation are based on rather classic video games like Super Mario Bros. [50], Angry Birds [97, 52] and Minecraft [59, 60]. Though using classical games makes it convenient for research, it can prevent AI-involved game design from being used in practice.
V-E Lower Barriers for Building New Games
There are some professional game engines to build simulators for machine learning such as Unity [80], Unreal Engine [^5] https://github.com/zfw1226/gym-unrealcv and Godot [160]. Developing a domain-specific platform for a game is not easy. Some simulators can only be used in specific scenarios, and if the AI algorithm needs to be used in other situations, a new simulator may need to be built, which is time-consuming. One way to make simulators more versatile is by applying description languages. VGDL [91] is a good example. Games written by VGDL can be compiled by a specific compiler such as GVGAI [71, 72, 92] and UnityVGDL [161]. Another way is using general platforms, especially professional game engines to build simulators such as Unity [80], Unreal Engine and Godot [160]. Using professional game engines, simulators can be built much easier and run faster. With remote procedure call, algorithms do not need to be implemented with the same programming language as simulators do, which makes it possible to use general platforms for AI research, reducing engineering issues.
V-F Large Language Models for Games
Recently, the emergence of LLMs such as ChatGPT [145] narrows the gap of game generation [121]. Using natural languages, i.e., interacting with LLMs, everyone can create their own game. Some other applications of LLMs for assisting game development such as level generation and dialogue generation also thrive. MarioGPT [20, 21] provides an end-to-end Mario level generator through a fine-tuned GPT2. Users can write some simple prompts such as “many pipes, some enemies” to generate game levels with their preference. The work of [19] validates the effectiveness of LLMs for generating Sokoban levels from the perspective of playability, novelty, diversity and accuracy.
Instead of assisting game content generation, LLM itself can also be regarded as a part of games. NPCs have been playing an essential role in games. Traditionally, NPCs have been confined to fixed scripts, limiting them to repetitive conversations that do not interact with or affect players’ strategies. LLMs, however, provide the potential of human-like NPCs, who may act and respond according to the behaviours of players. For example, Park et al. [17] adopt ChatGPT for the generative agent design. Players directly communicate with NPCs controlled by ChatGPT. Kumaran et al. [18] propose SceneCraft based on LLM for narrative generation in a 3D game. In addition to dialogue, SceneCraft generates emotion and gestures of NPCs during the interactions with human players. The work of [162] shows the possibility of developing software including games using multiple LLM agents.
Notably, LLMs also show the potential of playing games. Wu et al. [22] use GPT-4 to read a paper about game Spring, and then directly apply GPT-4 to play the game. According to [22], GPT-4 presents outstanding performance than advanced RL algorithms without training. It is widely believed that LLMs, with their remarkable capabilities and capacity for generalisation, are driving the future direction of game and game AI development.
V-G Roles of Games in Interdisciplinary Research in AI
Games not only play an important role in the development of AI, but also is an ideal tool for studying interdisciplinary research in AI and other computing-related areas, such as social intelligence of AI [163], affective computing [164], health and education [165, 166, 167], cognitive behavioural therapy for children [168]. Besides, game-based learning has presented its strong capability in developing problem-solving and cooperative skills of children, compared with traditional lessons [169]. Games have also been used as the mediums for educating AI [170].
VI Conclusion
This paper reviews and categorises the publicly available games and game-based platforms for AI research. Growing research interests in AI for game testing, and AI for game design, games for artificial general intelligence, games for creative AI, have been induced. We present the games and game-based platforms for playing according to their characteristics, such as aims, the number of agents, observability, programming languages, and whether AI related competitions are held. Games and platforms for design are grouped by focused content type, such as 2D levels, 3D levels, narratives, rhythm charts, and cards.
Games serve as controllable and customisable simulators. Consequently, we discuss existing challenges thriving from games. The rapid development of games leads to the curse of high dimensions. Poor exploration and partial observability are also commonly found in games since agents need to interact with the environment. Generalisation capability of AI agents is validated on games, which poses the question if they still work in some unseen games and levels. The multi-agent games, where non-stationarity, coordination and human-AI interaction are involved, are also discussed.
Nevertheless, this paper presents how AI works for games and games work for AI. As a perfect assistant to humans, AI is capable of testing games and balancing game content during the process of game development. Besides, AI can be applied to generate different types of game content, such as levels, music and stories, which highly releases human resources. Open-world games provide virtual worlds for artificial general intelligence to explore. As the reflection of real-world scenarios, games can be applied as preliminary test cases. Furthermore, the recent emergence of LLMs offers a new perspective on the development of AI techniques. Natural languages can be directly used for creative design, such as levels, images and the games themselves. Recent studies present how LLMs make actions in games according to the textual representations. LLMs also show the potential of directly engaging in gameplay as NPCs, which entertains human players.
However, some concerns about the development of game AI platforms are raised. Although AI works for game testing and design, it lacks a united framework across games. Essential efforts have to be put in when applying AI to testing new games or designing game content. Available interfaces for modern commercial games are limited by some engineering issues. A large number of commercial games do not provide official APIs for AI research. Although existing platforms offer a controllable, safe and diverse environment to study, explore, evaluate and experiment with different ideas, the augmentation of test-beds for AI would be facilitated by more accessible interfaces for building new games. Considering the gaps between academics and industries that own plenty of computational resources and copyright of commercial games [171], getting started for beginners in game AI on some simple environments such as Atari may be a good choice. It is expected to see new open-source game-based platforms with under-explored game genres, diverse content types, user-friendly APIs and documents.
Finally, new, open-ended questions arise while this paper was being written and, at the same time, LLMs are rapidly advancing almost with daily improvements: (i) What challenges that motivated the implementation of the games/platforms reviewed by this paper or were posed by those games/platforms cannot be solved by LLMs in the near future and remain valuable to study? (ii) What the games/platforms reviewed by this paper can contribute to the development of LLMs?

### v4

@@
-Game-based Platforms for Artificial Intelligence Research
+Games for Artificial Intelligence Research: A Review and Perspectives
 Abstract
@@
 Those platforms provide ideal benchmarks for exploring and comparing artificial intelligence ideas and techniques.
-This paper reviews the game-based platforms for artificial intelligence research, provides guidance on matching particular types of artificial intelligence with suitable games for testing and matching particular needs in games with suitable artificial intelligence techniques, discusses the research trend induced by the evolution of those platforms, and gives an outlook.
-Impact Statement:
+This paper reviews the games and game-based platforms for artificial intelligence research, provides guidance on matching particular types of artificial intelligence with suitable games for testing and matching particular needs in games with suitable artificial intelligence techniques, discusses the research trend induced by the evolution of those games and platforms, and gives an outlook.
+{IEEEImpStatement}
 Games have been playing an essential role in the development of AI techniques and education by serving as valuable test-beds.
@@
 Besides, the paper presents the evolution of games along with the advancement of AI techniques, thereby igniting inspiration for future investigations.
-Readers from industry will also discover how AI can impact games and how games can impact AI research through this survey.
-Index Terms:
-Game AI, platform, reinforcement learning, procedural content generation, competition
-I Introduction
+Readers from industry will also discover how AI can impact games and how games can impact AI research through this review.
+{IEEEkeywords}
+Game AI, computer games, reinforcement learning, procedural content generation, competition
+1 Introduction
 Figure 1: Roadmap of games and platforms, created by authors using screenshots of the games and platforms and figures of papers cited in this work.
+[…]
 Games have played an essential role in the rapid development of artificial intelligence (AI).
-As a controllable, cost-effective, and realistic simulator [1], games provide a virtual world for AI to explore.
-A variety of real-life problems, such as real-time decision making, scheduling, design and education, can be reflected in such virtual worlds.
-Games have naturally become the perfect test-beds to verify and explore the ability, generality, robustness and safety of AI.
-Game-based platforms have provided an ideal playground for AI researchers to study, explore, evaluate and experiment with different ideas in a controllable and safe environment.
-Such platforms also enable us to study and answer what-if questions that are hard to do safely in the real world.
+As a controllable, cost-effective, and realistic simulator [1], games provide virtual worlds for AI to explore.
+A variety of real-life problems, such as real-time decision making [2, 3], scheduling [4], design [5, 6] and education [7], can be reflected in such virtual worlds.
+Games have naturally become the perfect test-beds to verify and explore the ability, generality, robustness and safety of AI [8, 9, 10].
+Games or game-based platforms have provided an ideal playground for AI researchers to study, explore, evaluate and experiment with different ideas in a controllable and safe environment [5, 11].
+Such games and platforms also enable us to study and answer what-if questions that are hard to do safely in the real world.
 They can even stimulate researchers to ask new research questions.
-Various games have been used in game-based platforms, from single-player to multi-player, and from puzzles to real-time strategy games, with the emergence of diverse game genres and mechanisms.
+Various games have been used in game-based research platforms, from single-player to multi-player, and from puzzles to real-time strategy (RTS) games, with the emergence of diverse game genres and mechanisms.
 Such diversity produces a large number of diverse challenges for AI to solve.
-For instance, it has been demonstrated in various games (e.g., Chess [2], Atari [3], Poker [4], Go [5], StarCraft II [6], Dota 2 [7], and Texas hold’em [8]) that AI can perform at a level beyond that of humans.
+For instance, it has been demonstrated in various games (e.g., Chess [12], Atari [2], Poker [13], Go [14], StarCraft II [15], Dota 2 [16], DouDizhu [17], Texas hold’em [18] and GuanDan [19]) that AI can perform at a level beyond that of humans.
 Besides playing games by AI, designing new games by AI is an emerging research area which helps exploit AI’s potential in creativity [20].
 Games are rich in their content types, including visuals, audio, levels, rules, etc., making them challenging and diverse environments for innovative AI research [5, 21].
-In terms of creative game design, search-based methods, machine learning (ML) and deep learning (DL) techniques have been used for procedural content generation (PCG) where both AI creativity and human-AI co-creativity are explored [12, 13, 14, 15, 16].
+In terms of creative game design, search-based methods, machine learning (ML) and deep learning (DL) techniques have been used for procedural content generation (PCG) where both AI creativity and human-AI co-creativity are explored [6, 22, 23, 24, 25, 26, 27].
 Emergent abilities of large language models (LLMs) also open up new opportunities and paradigms of playing and designing games [28, 29, 30, 31, 32, 33].
 More recently, Bruce et al.
-from DeepMind introduce Genie, the generative interactive environment that is capable of learning from videos to generate playable, endless 2D platformer games from an image prompt [23].
-AI-generated contents can help enhance the generality of AI itself [11].
-AI techniques can help to automate the design tasks to accelerate the process, generate novel and diverse contents, and personalise contents for individual players.
-This paper provides a bird’s-eye view of publicly available game-based platforms or games for AI research, aiming to provide a brief but comprehensive overview for researchers and students that want to either use AI for games or use games for AI.
+[34] from DeepMind introduce Genie, the generative interactive environment that is capable of learning from videos to generate playable, endless 2D platformer games from an image prompt.
+AI-generated contents can help to enhance the generality of AI itself [21].
+AI techniques can help to automate the design tasks to accelerate the process, generate novel and diverse contents, and personalise contents for individual players [6, 26].
+This paper provides a bird’s-eye view of publicly available game-based platforms and games for AI research, aiming to provide a brief but comprehensive overview for researchers and students that want to either use AI for games or use games for AI.
 Fig.
@@
 The paper is organised as follows.
-Section II explains the scope and methodology of this survey.
-Taxonomies of playing and designing tasks involving games for AI research are provided in Sections III and IV, respectively.
-The research trends and challenges induced by the emergence and evolution of those game-based platforms are discussed in Section V.
-Finally, Section VI concludes.
-II Methodology of Survey
+Section 2 explains the scope and methodology of this survey.
+Taxonomies of playing and designing tasks involving games for AI research are provided in Sections 3 and 4, respectively.
+The research trends and challenges induced by the emergence and evolution of those games and game-based platforms are discussed in Section 5.
+Finally, Section 6 concludes.
+2 Methodology of Survey
 Our survey considers the interpretation of games as Liu et al.
 [6], “any games a human would conceivably play, including board games, card games, and any type of video games, such as arcade games, role-playing games, first-person shooters, puzzle games, and many others”.
-Simulators for autonomous driving or robotic control are not covered since they are specifically designed for simulating and studying robotic systems with accurate physics simulations.
-To simplify the language, from now on, we use the term “game” and “platform” to refer to a game simulator and an environment with multiple games for AI research, respectively.
-Yannakakis and Togelius [10] have enumerated and discussed some game-based platforms for game playing and design till 2018.
+To simplify the language, from now on, the term “game” refers to either a game simulator or an environment with multiple games, while “game-based platforms” (or “platform” in short) specially points to the latter.
+Yannakakis and Togelius [5] have enumerated and discussed some games for playing and design till 2017.
 Shao et al.
-[24] have reviewed some video game platforms for researching deep reinforcement learning till 2019.
+[35] have reviewed some video games for researching deep reinforcement learning (deep RL), which were released before 2019.
 The work of Swiechowsk [1], Giannakos et al.
-[25], Duan et al.
+[36] and Duan et al.
 [37] have also touched on this topic.
-However, to our best knowledge, no existing paper has provided a complete review or taxonomy on recent game-based platforms, while more than 56 new ones have been released since 2018.
+However, to our best knowledge, no existing paper has provided a complete review or taxonomy on recent games.
+As a side note, the fields of game AI, autonomous driving,robotics and etc., share some similar challenges [38, 39].
+On one hand, some of the games, simulators or virtual environments may be used by one or more fields.
+On the other hand, the AI approaches and techniques widely used in one field can be applied to another.
+However, to be more focused in this survey, virtual environments specifically designed for simulating and studying autonomous driving and robotic systems, such as Carla [40], Omniverse [41] and Mujoco [42] with accurate physics simulations, are not covered.
 The following survey methodology is applied.
-First, we have gone through the competitions organised at related conferences in AI or game AI, including NeurIPS/NIPS, IJCAI, AAAI, ICML, ECML, ACML, AIIDE, IEEE CIG, IEEE CoG, ACM FDG, ACG, GECCO, WCCI/CEC, between 2016 and February 2024, and collected corresponding platforms.
-The year 2016 is chosen as AlphaGo [27] has boosted the research interest in game AI.
+First, we have gone through the competitions organised at related conferences in AI or game AI, including NeurIPS/NIPS, IJCAI, AAAI, ICML, ECML, ACML, AIIDE, IEEE CIG, IEEE CoG, ACM FDG, ACG, GECCO, WCCI/CEC, between 2016 and February 2024, and collected corresponding games/ platforms.
 Then, we searched with Google Scholar and Web of Science using the following search terms, “game” AND (“artificial intelligence” OR “AI”) AND (“design" OR “generation”) AND (“platform” OR “interface” OR “application” OR “software”), for related papers published or available online between 2016 and February 2024, assuming that most of the earlier ones can be found in Yannakakis and Togelius [5], Shao et al.
@@
 Finally, for all related papers that did not specify the link to the corresponding platform or code, we searched for the code on the authors’ personal pages, in GitHub, and using Google Search.
-Again, only publicly available ones are included in our survey.
+The year 2016 is chosen for the following reasons.
+AlphaGo [43] introduced in 2016 is a milestone for AI research.
+Its success presents the potential of several AI approaches and techniques, such as deep RL and search methods, in mastering complex games with vast search spaces and strategic depth beyond the human level, and inspires not only research on game AI [44, 45] but also interdisciplinary research, such as AI for science [46, 47] and AI for social interaction [48].
+As discussed previously, most of the related games published before 2017 have been covered by [5, 35].
+However, lots of new games have emerged as shown in this survey.
+Again, only publicly available ones are included.
 It’s worth mentioning that a platform or game may have been available online before the corresponding paper is published.
@@
 Such games are also included.
-III Playing Games by AI
+3 Playing Games by AI
 Games considered in this survey can be briefly categorised into tabletop games and video games.
@@
 This poses a challenge of extracting high-dimensional but uninformative states and action features.
-Moreover, credit assignment in games may be challenging due to the game length, unobservable behaviours of players, combinational actions and sparse rewards.
+Moreover, credit assignment in games may be challenging due to the game length, hidden behaviours of players, combinatorial actions and sparse rewards.
 Tab.
-I, Tab.
-II and Fig.
-3 categorise the games according to the game genre and the simulated physical space with which avatar(s) interact.
-Platforms equipped with multiple games are also introduced.
-TABLE I: 36 games for playing by AI.
+1, Tab.
+2 and Fig.
+3 categorise the games and platforms according to the game genre, research aims, the number of agents and observability of the simulated physical space with which agent(s) interact, as well as programming language and if they have been used by any AI-related competitions.
+Table 1:  40 games for playing by AI.
 The column with header C indicates whether an AI competition is associated.
@@
 “1”, “2” and “>2” denote the number of agents, representing single-, two- and multi-agent.
-, and denote whether the platform is deterministic, stochastic or both of them.
-$*{}^{*}$ highlights commercial game with publicly available API.
-“A”, “C”, and “M” denote adversarial, cooperative, competing as mixed scenarios.
+, and denote whether the platform is deterministic, stochastic or both.
+highlights commercial games with publicly available API.
 Type Platform Aim # Agents Observability Language Competition Reference
 Planning Learning $1$ $2$ $>2$ Full Partial
-Tabletop Games (Section III-A) Board ColorShapeLinks A C# [28]
-Blood bowl A Python [29]
+Tabletop Games (Section 3.1) Board ColorShapeLinks A C# [49]
 Mjai A JavaScript [50]
 Mjx A Python [51]
+Blood bowl A Python [52]
 Diplomacy M Python [53]
-Card ben M Python [33]
+Card VGC A Python [54]
+Tales of Tributes A C# [55]
 Sabberstone A C# [56]
-VGC A Python [35]
-Tales of Tributes A C# [36]
+Fireplace A Python [57]
+ben M Python [58]
 Hanabi A M Python [59]
@@
 Text Light M M Python [62]
-Video Games (Section III-B) Grid- based 2D SimCity Python [41]
+Video Games (Section 3.2) Grid- based 2D Snake AI A Java [63]
+SimCity Python [64]
 Nethack Python [65]
-Snake AI A Java [43]
+Pommerman M Java/Python [66]
+Neural MMO M Python [67]
 microRTS A Java [68]
 Tribe A A Java [69]
-Overcooked C Python [46]
+Overcooked! C Python [70]
 Cuisine World C C++ [71]
-Pommerman M Java/Python [48]
-Neural MMO M Python [49]
 Pixel- based 2D Mario AI Java [72]
 MsPacMan-vs-Ghosts Java [73]
-AIBIRDS Java/Python [52]
-XFC Python [53]
 Geometry Friends C C# [74]
 AIGD2 A Java [75]
+XFC Python [76]
 FightingICE A Python [77]
-Honor of King $*{}^{*}$ A Python [57]
-Dota 2 $*{}^{*}$ M Python/ Lua [7]
-StarCraft II $*{}^{*}$ A M C++/Python [6]
-3D ViZDoom Python [58]
-MineRL Python [59]
+AIBIRDS Java/Python [78]
+nsh Python/Lua [79]
+Honor of King  A Python [80]
+Dota 2  M Python/Lua [16]
+Dunk City Dynasty  M Python [81]
+StarCraft II  A M C++/Python [15]
+3D MineRL Python [82]
 MineDojo Python [83]
+Google Research Football M Python [84]
+ViZDoom Python [85]
 Hide and Seek M Python [86]
- Single-player deterministic game  Single-player stochastic game Configurable    $*{}^{*}$ Commercial game with API
+ Single-player deterministic game  Single-player stochastic game Configurable Commercial game with API
 Games with #Agent2absent2$\geq$ 2≥ 2: “A” for adversarial “C” for cooperative “M” for mixed
-TABLE II: 20 platforms with multiple games for playing by AI (cf.
+Table 2: 20 platforms with multiple games for playing by AI (cf.
 caption of Tab.
-I for reading instructions).
+1 for reading instructions).
 Whether a game is adversarial or collaborative is not indicated as each platform contains a number of different games.
@@
 “1”, “2” and “>2” denote the number of agents, representing single-, two- and multi-agent.
-Action-adventure game (AAG) involves video games such as puzzle and combat games.
+Action-adventure game (AAG) involves video games such as puzzles and combat games.
 Platform Aim Game genre #Agents Observability Language Competition Reference
@@
 OpenSpiel C++/Python [89]
+TAG Java [90]
 Ludii Java/Python [91]
-TAG Java [66]
 Pgx Python [92]
@@
 PettingZoo Python [94]
-Minigrid Java/Python [70]
 GVGAI Java/Python [95]
 GVGAI gym Java/Python [96]
+Minigrid Java/Python [97]
+Melting Pot Python/Lua [98]
 Stratega C++/Python [99]
 Griddly Python [100]
-Melting Pot Python/ Lua [75]
 Retro Python [101]
@@
 DeepMind Lab Python [104]
+MiniWorld Python [105]
 ML-agents Python [106]
-MiniWorld Python [81]
-III-A Tabletop Games
+3.1 Tabletop Games
 Board, card, puzzle games and other tabletop games raise several challenges to AI, such as partial observability and multi-agent cooperation or competition.
 In ColorShapeLinks [49], players win by connecting pieces with the same colour or shape.
-The platform of ColorShapeLinks not only provide a visual front-end via the Unity game engine, but also a pure text-based one.
-ColorShapeLinks has been used for education and held as an AI competition, thus it can be easily accessed by individuals with different backgrounds.
+The platform of ColorShapeLinks provide not only a visual front-end via the Unity game engine, but also a pure text-based one.
+ColorShapeLinks has been used for education and a testbed of an AI competition.
+It can be easily accessed by individuals with different backgrounds.
 However, ColorShapeLinks is relatively simple for research purposes since good solutions can be quickly found.
+Mjai [50] and Mjx [51] are partial-observable AI environments of riichi Mahjong.
+Thirteen tiles on the own hand and discarded cards are available for the player.
+If a special permutation of cards is formed, the player wins the game.
+Mjx is 100 times faster than Mjai in running time and is compatible with Gym’s API [51].
 Blood bowl [52] simulates a fully-observable board game that two players control their own teams and are rewarded with touchdowns.
@@
 In Blood bowl, the observation space consists of both spatial and non-spatial information and the action space differs across states, which leads to a new challenge in representation.
-Mjai [30] and Mjx [31] are partial-observable AI environments of riichi Mahjong.
-Thirteen tiles on the own hand and discarded cards are available for the player.
-If a special permutation of cards is formed, the player wins the game.
-Mjx is 100 times faster than Mjai in running time and is compatible with Gym’s API [31].
 Diplomacy is a seven-person board game, considering both cooperation and competition [53].
 In this game, AI agents are required to understand human players’ imitations and behaviours.
-Hearthstone: Heroes Of Warcraft is a popular commercial turn-based collectable card game (CCG).
-An open-source implementation version called Sabberstone [34] is released for playing and deck construction.
 Pokémon Video Game Championships (VGC) is a turn-based strategy game [54].
@@
 During the battles, player agents decide the order of Pokémons and specify the skills to attack the opponents according to previous information.
-The work of Budzki et al.
-[36] introduces a deck-building card game, Tales of Tribute.
+The work of [55] introduces a deck-building card game, Tales of Tribute.
 Players build the deck with the same initial cards during the game through battles and trades, meeting some dynamics in a long-term strategy.
 Tales of Tribute also provides a graphic interface and supports implementing AI agents with C#.
-The three games are all CCG, where players build decks and fight against opponents, but they have different research purposes.
-Sabberstone and Tales of Tribute both focus on the decision agents, i.e., how to play cards.
-However, Tales of Tribute also considers building decks, while Sabberstone doesn’t.
+Hearthstone: Heroes Of Warcraft is a popular commercial turn-based CCG.
+There are open-source implementations, e.g., Sabberstone [56] and Firepalce [57], for playing.
+The three games are all collectable card games (CCG), where players build decks and fight against opponents, but they have different research purposes.
+Sabberstone, Firepalce and Tales of Tribute focus on researching decision-making agents, i.e., how to play cards.
+However, Tales of Tribute can also be used to study building decks by AI.
 Meanwhile, VGC involves meta-game balance that adjusts attributes of cards during the battles.
+Bridge is another challenging board game for AI because of its partial observability, high branching factor, cooperation and competition [107].
+23 editions of the World Computer-Bridge Championship have been organised with the latest one in 2019222World Computer-Bridge Championship: https://bridgebotchampionship.com/.
+The game engine of the championship was not found by authors.
+However, there is a recently released game engine called ben [58] for AI research in Bridge.
 Different from some traditional card games, players of Hanabi [59] can only see the cards of other players.
@@
 As a fully cooperative game, players need to form fireworks that consist of the same colour cards in ascending order to win.
-Besides Poker [4] and Texas hold’em [8], Bridge is another challenging board game for AI because of its partial observability, high branching factor, cooperation and competition involved in the game [82].
-23 editions of the World Computer-Bridge Championship have been organised with the latest one in 2019222World Computer-Bridge Championship: https://bridgebotchampionship.com/.
-The game engine of the championship was not found by authors.
-However, there is a recently released game engine called ben [33] for AI research in Bridge.
-In most games, agents directly control the avatars to move or play cards, while in some others, agents take non-direct actions such as languages.
+In most games, agents directly control avatars to move or play cards, while in some others, agents take non-direct actions such as languages.
 The social card game, The Chef’s Hat [60] is also implemented based on Gym [93] and used for RL research.
@@
 Light is a text-based adventure game [62], with which dialogue-based human-AI interaction is studied.
-Tabletop games usually don’t require specific designs in state representation, while the action space in those games changes during an episode.
-For example, in Go, the state is easily constructed by a matrix, where three symbols or numbers denote white, black or empty.
+Usually, no specific design is required for representing states in tabletop games, while the action space in those games changes during an episode.
+For example, in the game of Go, the game state is easily constructed by a matrix, where three symbols or digits denote white, black or empty.
 However, the empty space for placing stones changes in a decreasing trend, leading to a variable action space.
 On the other hand, tabletop games often involve multiple players with complex and even dynamic relationship (collaboration vs.
-competition), pose new challenges and encourage studies in various research fields such as multi-agent systems [83], game theory [84] and psychology [85].
-III-B Video Games
-This section categorises the video games into grid-based 2D, pixel-based 2D and 3D according to the simulated physical space with which avatar(s) interact.
+competition), pose new challenges and encourage studies in various research fields such as multi-agent RL [108, 109, 110], game theory [111, 112] and psychology [113].
+3.2 Video Games
+This section categorises video games into grid-based 2D, pixel-based 2D and 3D according to the simulated physical space with which avatar(s) interact.
 Grid-based games are usually constructed by some preset tiles where players act with a set of discrete actions.
@@
 Furthermore, the difference between games and 3D virtual environments for robotics and autonomous driving, such as MuJoCo [42] and CARLA [40], is narrowing.
-III-B1 Grid-based 2D Games
+3.2.1 Grid-based 2D Games
 Classic single-player and multi-player competitive games have been implemented using grid-based maps.
-SimCity [41] is an open-ended city-building 2D video game with an open-source version provided for reinforcement learning research.
+Snake AI [63] allows two snakes to fight for an apple.
+Since the apple is placed randomly in maps, the game can sometimes be unbalanced.
+SimCity [64] is an open-ended city-building 2D video game with an open-source version provided for RL research.
 Nethack [65] is a single-agent roguelike game with procedurally generated levels.
-Agents meet the different stages at each episode, which drives AI’s abilities for long-term exploration and generalisation.
+Agents meet the different scenarios at each episode, which drives AI’s abilities for long-term exploration and generalisation.
 Compared with existing single-player games, Nethack provides a faster simulator with massive tasks and game dynamics.
 Multi-agent games involve multiple independent decision agents, oriented towards cooperative and competitive behaviours.
-Snake AI [43] allows two snakes to fight for an apple.
-Since the apple is placed randomly in maps, the game can sometimes be unbalanced.
+Pommerman [66] is a multi-player competitive game for at least four players, who start with a limited number of bombs.
+Players move in a grid-based map and plant the bombs to destroy their opponents.
+Pommerman allows multiple players to cooperate and compete as a team on the battlefield.
+Neural Massive Multiplayer Online (Neural MMO) [67] considers a large multi-agent scenario where at most 1024 agents learn to cooperate and compete to survive in the grid-based world at the same time.
 Modern strategy games incorporate digital elements to achieve complex and diverse mechanisms.
-microRTS [44] is a lightweight, grid-based real-time strategy (RTS) game framework that has fewer rendering sprites and focuses on the game’s mechanics, in which units are represented by different colours and shapes.
+microRTS [68] is a lightweight, grid-based RTS game framework that has fewer rendering sprites and focuses on the game’s mechanics, in which units are represented by different colours and shapes.
 Compared with StarCraft II, microRTS is easier to manipulate, so that researchers can focus more on the design of AI ideas and algorithms.
@@
 Compared to the implementation by [70], CuisineWorld is more suitable for research on LLMs.
-Pommerman [48] is a multi-player competitive game for at least four players, who start with a limited number of bombs.
-Players move in a grid-based map and plant the bombs to destroy their opponents.
-Pommerman allows multiple players to cooperate and compete as a team on the battlefield.
-Neural Massive Multiplayer Online (Neural MMO) [49] considers a large multi-agent scenario where at most 1024 agents learn to cooperate and compete to survive in the grid-based world at the same time.
 Grid-based games usually have simplified state and action space, as well as clear game mechanisms, which allow to train and test AI agents at a faster speed.
 However, this simplicity limits its usage in AI approaches considering continuous decision space and complex strategies.
-III-B2 Pixel-based 2D Games
+3.2.2 Pixel-based 2D Games
 Pixel-based games support more complex game mechanisms with continuous state space, while new challenges are introduced.
 Mario AI framework [72] is based on the game Super Mario Bros., where players control Mario to run and jump through some barriers in a partially observable environment to arrive at the endpoint.
-However, the provided levels are simple and a limited number of levels are available, planning algorithms such as $*{}^{*}$ can quickly solve it without leveraging sophisticated AI approaches.
+However, the provided levels are simple and a limited number of levels are available, planning algorithms such as A can quickly solve it without leveraging sophisticated AI approaches.
 MsPacMan-vs-Ghosts [73] is the implementation of Ms.
@@
 Both full and partial observability are supported in MsPacMan-vs-Ghosts.
+Geometry Friends [74] provides a cooperative scenario that two agents aim to solve puzzles together, restricted by real physics conditions such as friction and gravity.
+A simple implementation of Planet Wars, named AIGD2 [75], brings a new challenge with extra-long skill-depth to RTS games.
+XFC [76] simulates the game Asteroid Smashers, in which AI controls a spacecraft to avoid asteroids and destroy pesky rocks.
+XFC designs a specific agent architecture, which is suitable for research on explainable AI and fuzzy AI.
+FightingICE [77] considers a 2D fighting video game that AI agents fight against opponents according to sound.
+“Blind" agents face challenges such as partial observability and insufficient training.
 AIBIRDS [78] is the platform with the popular puzzle game Angry Birds.
 Agents are required to shoot pigs to destroy the building made of blocks of different types and shapes.
-XFC [53] simulates the game Asteroid Smashers, in which AI controls a spacecraft to avoid asteroids and destroy pesky rocks.
-XFC designs a specific agent architecture, which is suitable for research on explainable AI and fuzzy AI.
-Geometry Friends [54] provides a cooperative scenario that two agents aim to solve puzzles, restricted by real physics conditions such as friction and gravity.
-A simple implementation of Planet Wars, named AIGD2 [55], brings a new challenge with extra-long skill-depth to RTS games.
-FightingICE [56] considers a 2D video fighting game that AI agents fight against opponents according to sound.
-“Blind" agents face challenges such as partial observability and insufficient training.
-Honor of Kings Arena, Dota 2 and StarCraft II [6] are three popular commercial games with publicly available APIs that can be used for AI research.
+nsh [79] is a fight game simulator for the commercial massive multiplayer online (MMO) game Justice Online.
+Renders and memory usage are largely reduced in nsh, while game rules are close to the ones in the original game.
 Honor of Kings Arena [80] provides an interface to interact with Honor of Kings 1v1, in which agents learn to fight against diverse opponents.
@@
 2 presents the heroes controlled by OpenAI Five [16].
+Dunk City Dynasty simulates a 3v3 basketball commercial game with the same name [81].
+A dataset with 400GB human data and records is also provided.
 An interface connecting to StarCraft II [15] is also released for research in multi-agent RL and map design.
@@
 Figure 3: Games and platforms for playing, categorised as board/card, grid-based 2D, pixel-based 2D, 3D, social and text game types.
-III-B3 3D Video Games
+3.2.3 3D Video Games
 3D video games provide a more realistic simulation of the real world.
 Decision agents need to take actions based on the 3D vision instead of a flat plane in 2D games, which requires effectively processing high-dimensional visual information to make effective decisions.
+MineRL [82] is based on Minecraft, where a large-scale human dataset is provided.
+MineDojo [83] provides a Gym-like environment based on Minecraft.
+Agents are supported to directly access the game for multiple tasks including material mining, crafting and building.
+Besides, a massive multimodal dataset, captured from video websites, gameplays, wiki pages and Reddit posts is provided.
+Google Research Football [84] provides a real-world football challenge with a physical 3D engine, in which multiple tasks with variable difficulty are available.
 ViZDoom [85] is a highly customisable doom-based environment with partial observability.
@@
 Its integrated editor allows users to manually create diverse task suits.
-ViZDoom is suitable as the test-bed for solving visual RL problems.
-MineRL [59] is based on Minecraft, where a large scale human dataset is provided.
+ViZDoom is suitable as the test-bed for visual RL research.
 The work of Baker et al.
 [86] presents a specific platform for Hide and Seek, which can be used for training and understanding multi-agent cooperative and competitive AI.
-MineDojo [60] provides a Gym-like environment based on Minecraft.
-Agents are supported to directly access the game for multiple tasks including material mining, crafting and building.
-Besides, a massive multimodal dataset, captured from video websites, gameplays, wiki pages and Reddit posts is provided.
-Figure 4: Summary of games and platforms used for studing AI for game design, created by authors using screenshots of the games and platforms, as well as figures from their corresponding papers, reviewed in Section IV.
-III-C Platforms
+Figure 4: Summary of games and platforms used for studing AI for game design, created by authors using screenshots of the games and platforms, as well as figures from their corresponding papers, reviewed in Section 4.
+3.3 Platforms
 Classic and modern board, card and puzzle games are often integrated into a single platform as the AI test-bed.
@@
 Besides some popular classic games, OpenSpiel [89] provides Poker and Bridge, as well as some classic algorithms and visual tools.
+Different from GGP and Botzone, Tabletop Games Framework (TAG) [90] is designed to study general AI in modern tabletop games and has implemented more than 20 games, including Uno, Love Letter and Virus!.
+An extra analysis tool is equipped in TAG for studying action space, branching factors, hidden information, and so on.
 Ludii [114] is another general game-playing system based on similar principles to Ludi [115] but with different mechanisms.
@@
 The description language is easy to read and understand according to [114].
-Different from GGP and Botzone, the Tabletop Games Framework (TAG) [66] is designed to study general AI in modern tabletop games and has implemented more than 20 games, including Uno, Love Letter and Virus!.
-An extra analysis tool is equipped in TAG for studying action space, branching factors, hidden information, and so on.
-Pgx [67] is a environment of board games for RL research.
-It is written in JAX and can do parallel simulation via GPU/TPU.
+Pgx [92] is an environment of board games for RL research.
+It is written in JAX and can do parallel simulations via GPU/TPU.
 Pgx can hold single-/multi-agent board games and Atari-like environments including Shogi, Go, Chess, and Backgammon.
 According to [92], Pgx can perform 10-100 times higher throughput compared to other Python libraries.
-OpenAI Gym [68] and PettingZoo [69] offer a number of grid-based and pixel-based 2D games.
-OpenAI Gym [68] provides an API standard for RL research and contains massive games, such as Atari, as well as a wide range of third-party environments.
+OpenAI Gym [93] and PettingZoo [94] offer a number of grid- and pixel-based 2D games.
+OpenAI Gym [93] provides an API standard for RL research, and contains massive games (e.g., Atari), and a wide range of third-party environments.
 However, Gym typically focuses on the single-agent setting.
@@
 As providing the standard API, Gym and PettingZoo are now the most used and compatible platforms, which enhance the evaluation and reproduction in the RL community with a higher speed.
+General Video Games AI (GVGAI) [95] platform and GVGAI gym [96] offer a larger number of grid-based single- and two-player games, and have been used in their associated single-player planing, two-player planing, and single-player learning competitions.
+Users can create new games and levels in GVGAI in an intuitive way using Video Game Description Language (VGDL) [117].
 Minigrid [97] is a single-agent platform compatible with OpenAI Gym.
 The included games are simple in rendering content but can implement complex mechanisms.
-General Video Games AI (GVGAI) [71] platform and GVGAI gym [72] offer a larger number of grid-based single- and two-player games, and has been used in its single-player planing, two-player planing, and single-player learning competitions.
-Users can create new games and levels in GVGAI in an intuitive way using Video Game Description Language (VGDL) [91].
+Melting Pot [98] is a platform with more than 50 diverse games for evaluating the generalisation of multi-agent RL.
 Stratega [99] evaluates the generalisation of strategy agents with multiple games and levels with an $n$-player turn-based setting.
@@
 Griddly [100] particularly provides grid-based real-time strategy games with various numbers of players, with a faster running speed and lower memory than Minigrid and GVGAI.
-Melting Pot [75] is a platform with more than 50 diverse games for evaluating the generalisation of multi-agent RL.
 The Retro Learning Environment [101] extends Gym, by incorporating popular video games from Super Nintendo Entertainment and Sega Genesis.
@@
 Besides, these pixel-based game platforms usually have lower scalability in adding new games and tasks than grid-based platforms such as GVGAI.
-DeepMind Lab [79] acts as a 3D learning environment which is based on some open source software such as ioquake3.
+DeepMind Lab [104] acts as a 3D learning environment which is based on some open-source software such as ioquake3.
 It provides a set of challenging 3D navigation and puzzle-solving tasks for agents to learn.
-The primary purpose of DeepMind Lab is to act as a working platform for results in AI, especially deep reinforcement learning.
-ML-agents [80] is an open-source platform that allows to create new games in a simpler and more flexible manner with the help of the Unity3D toolkit.
-MiniWorld [81] provides a more accessible way to construct 3D environments for reinforcement learning and robotics research.
+The primary purpose of DeepMind Lab is to act as a working platform for results in AI, especially deep RL.
+MiniWorld [105] provides a more accessible way to construct 3D environments for RL and robotics research.
 It is qualified to generate and simulate environments with rooms, doors, mazes and other objects.
 MiniWorld is simpler than DeepMind Lab and ViZDoom in the view of engineering, so beginners can get started quickly without mastering Unity3D.
-III-D Guidance for Playing Games
+ML-agents [106] is an open-source platform that allows to create new games in a simpler and more flexible manner with the help of the Unity3D toolkit.
+3.4 Guidance for Playing Games
 According to the aim, number of agents, observability, language and competition, games and platforms reviewed in this work are categorised in Tabs.
-I and II.
+1 and 2.
 Those games and platforms are often developed for specific research purposes.
-Card/board-related games and platforms, including Blood bowl [29], Sabberstone [34] and Botzone [63], are suitable for research on tree search, causality and competitive scenarios in multi-agent systems.
-GVGAI [71, 72, 92], Stratega [73] and Procgen [77] are developed for the purpose of evaluating the generalisation of AI agents.
-Both GVGAI and Stratega can be used for research on planning and learning, while Procgen specially focus on RL.
+Readers can choose suitable games and platforms according to the categories.
+For example, if one focuses on learning-based research like RL [118] and imitation learning [119], games and platforms with ✓under Learning in the column Aim of Tabs.
+1 and 2 may be good choices.
+Examples of related research using those platforms include the work of Mnih el al.
+[2] using Atari, and Vinyals el al.
+[15] using StarCraft II.
+For research on searching algorithms such as Monte-Carlo tree search [120, 121], games and platforms marked in Planning column are good choices.
+With respect to different scenarios considering the number of agents, column #Agents provides such information.
+Single-agent scenarios usually meet deterministic, stochastic and configurable environments, denoted with ,   and , respectively.
+Sparse reward [122, 123] and long-term exploration [124] are common research topics involved with these scenarios.
+Besides, three multi-agent cases including adversarial, cooperative and mixed scenarios are denoted by “A", “C" and “M", respectively.
+Research on game theory [111, 112] and multi-agent RL [108, 109, 110] can benefit from the information for choosing suitable testbeds for experiments.
+Observability of games and platforms can also be checked in Tabs.
+1 and 2, which helps with research in partially observable planning [125] and RL [126].
+Card/board-related games and platforms, including Blood bowl [52], Sabberstone [56] and Botzone [88], are suitable for research on search, causality and competitive scenarios in multi-agent systems.
+GVGAI [127], Stratega [99] and Procgen [102] are developed for the purpose of evaluating the generalisation of AI agents.
+Both GVGAI and Stratega can be used for research on planning and learning, while Procgen specially focuses on RL.
 In the single-agent setting, Nethack [65] considers how an AI agent solves exploration, planning, and inventory management problems in a procedurally generated 2D environment.
@@
 However, game levels in Procgen can only be generated randomly.
-It is recommended to get started from Gym [68] and PettingZoo [69] for single- and multi-agent RL research since they are the most common platforms.
+It is recommended to get started from Gym [93] and PettingZoo [94] for single- and multi-agent RL research since they are the most commonly used platforms.
 Some platforms such as DeepMind Lab [104] and ML-agent [106] usually are equipped with some visual tools and loggers to analyse the behaviour of agents, but the requirement to learn about the game engines like Unity is crucial.
 From the two tables, Python is the most used programming language, which aligns with the trend of the famous Python libraries such as Torch and TensorFlow.
-IV Designing Games by AI
-TABLE III: 25 environments for designing games by AI.
+4 Designing Games by AI
+Table 3: 26 environments for designing games by AI.
 Although Mario-AI, HOG2, PhyDSL, StepMania, OSU!
-and MiniDungeons are available before 2016, they have been used for testing and evaluating the procedurally generated SMB levels in recent years.
+and MiniDungeons are available before 2016, they have been used for testing and evaluating the procedurally generated Super Mario Bro levels in recent years.
 Type Environment Language Competition Reference
-Game Tuning (Section IV-A) Parameter FightingICE Java [93]
+Game Tuning (Section 4.1) Parameter FightingICE Java [128]
 AIGD2 Java [75]
 VGC Python [54]
-Content Generation (Section IV-B) 2D level Mario AI Java [50]
+RaidEnv Python [129]
+2D level Science Birds Java [130]
+Content Generation (Section 4.2) Mario AI Java [72]
+Morai-Maker-Engine C# [131]
 HOG2 C++ [132]
@@
 Talakat JavaScript [134]
-Science Birds Java [97]
-Morai-Maker-Engine C# [98]
+Keiki Python [135]
 PCGRL gym Python [136]
 micro-rct Python [137]
-Keiki Python [101]
 3D level GDMC Python [138]
@@
 Metastone Java [145]
-Other (Section IV-C) Code PhyDSL Java [110]
+Other (Section 4.3) Code PhyDSL Java [146]
 Persona MiniDungeons Java [147]
@@
 Tab.
-III and Fig.
-4 summarise the games and platforms used for game design.
+3 and Fig.
+4 categorise the games and platforms used for game design according to research aim including game tuning and content generation, manipulated content type (e.g., parameter, 2D level, 3D level and narrative), programming language and if they have been used by any AI-related competitions.
+Some resources such as code and dataset are labelled as the other type.
 Those platforms provide diverse and representative test-beds for discovering the potentials in AI-involved game design.
-IV-A Game Tuning
+4.1 Game Tuning
 Game tuning is vital since it directly affects players’ experience.
 Unsuitable game parameter settings may lead to game unbalance or low quality, where players find it hard to play the game or feel bored.
-Generally, all the games introduced in Section III that involve configurable parameters can be tuned with search-based methods for different purposes, such as adjusting the game difficulty.
+Generally, all the games introduced in Section 3 that involve configurable parameters can be tuned with search-based methods for different purposes, such as adjusting the game difficulty.
 For instance, FightingICE [77] has been used to study the dynamic difficulty and battle sound adjustment [151].
+AIGD2 [75], an open-source game similar to Planet Wars, allows to explore the parameter space, which can significantly affect the gameplay via GVGAI agents [152].
 VGC [54] allows adjusting the attributes of Pokémons, called meta-game balance.
 Compared to other platforms, it specifically facilitates the investigation of enabling fair competition in gameplay.
-AIGD2 [55], an open-source game similar to Planet Wars, allows to explore the parameter space, which can significantly affect the gameplay via GVGAI agents [116].
-IV-B Content Generation
+RaidEnv [129] aims to optimise boss raid scenarios.
+Skills and statistics like health and attack can be adjusted by AI techniques, meeting the needs of different situations.
+4.2 Content Generation
 Content generation is involved with applying AI techniques to create new game content.
-The generated content types include 2D level, 3D level, narrative, rhythm chart and card, following Liu et al.
+The generated content types include 2D levels, 3D levels, narratives, rhythm charts and cards, following Liu et al.
 [6].
 Different content types are represented in different data structures, raising various challenges to content generation research.
-IV-B1 2D Map & Level
+4.2.1 2D Map & Level
 Game level is the most commonly generated content type.
 A 2D level is usually presented in a plane, where agents move in two dimensions.
+Science Birds [130] provides a generative platform, which can load XML-based level files to play Angry Birds, a physics-based puzzle video game, which requires the level components’ physical feasibility and stability.
+Generating levels for physical-based puzzle games with AI is complex since more physical constraints, larger state space and action space should be considered.
+Angry Birds AI Level Generation Competition [78] uses Science Birds as its simulator.
 Mario-AI-Framework [72], a fully-Java copy of Super Mario Bros., provides a tile-based representation of game levels.
@@
 5 shows examples of Mario levels generated by [153], tested and visualised in Mario-AI-Framework [72].
-Hierarchical Open Graph 2 (HOG2) [94] includes several grid-based puzzle games such as Snakebird [118] for exhaustive PCG [119].
-Morai-Maker [98] is a Unity3D-based platform for Super Mario Bros., which provides three built-in machine learning-based 2D level generators.
+Morai-Maker [131] is a Unity3D-based platform for Super Mario Bros., which provides three built-in ML-based 2D level generators.
 Co-creation between humans and AI is supported by the platform with a visible game level editor [131], which distinguishes Morai-Maker from other Super Mario Bros.
 platforms.
+Hierarchical Open Graph 2 (HOG2) [132] includes several grid-based puzzle games such as Snakebird [154] for exhaustive PCG [155].
 Procedural Zelda [133] is a grid-based game platform for level, combat and puzzle generation.
 Both offline and online control are supported during the generation with data logging capabilities.
-PCGRL gym [99] implements RL environments for grid-based level generation with the Gym interface.
-Binary, Zelda and Sokoban problems are integrated into PCGRL gym.
+Talakat [134] is a bullet hell game engine which defines a domain-specific language for bullet hell levels, supporting the grammar evolution of bullet hell game content.
+Keiki [135] is another bullet hell type game engine which provides a native Python interface describing bullet hell barrage.
+While Talakat represents barrages as scripts, Keiki provides a function to convert barrage scripts into vector-based representations, making it convenient for ML-based content generators.
+Procedural content generation via reinforcement learning (PCGRL) gym [136] implements RL environments for grid-based level generation with the Gym interface.
+Binary, Zelda and Sokoban are integrated into PCGRL gym.
 micro-rct [137] is a theme park management simulator, inspired by RollerCoaster Tycoon 2.
-In [100], researchers use a controllable PCGRL to generate game map.
+In Earle et al.
+[137], researchers use a controllable PCGRL to generate game maps.
 micro-rct is used to evaluate the controllability of PCGRL.
 Unlike other platforms which serve for search-based, supervised learning-based, or unsupervised learning-based content generation algorithms, PCGRL gym [137] provides a standardised benchmark for RL-based PCG, however, no accessible playing interface for human players is provided.
-Science Birds [97] provides a generative platform, which can load XML-based level files to play Angry Birds, a physics-based puzzle video game, which requires the level components’ physical feasibility and stability.
-Generating levels for physical-based puzzle games with AI is complex since more physical constraints, larger state space and action space should be considered.
-Angry Birds AI Level Generation Competition [52] uses Science Birds as its simulator.
 Figure 5: Mario level segments generated by the work of [153] tested and visualised in Mario AI framework [72].
@@
 Meanwhile, the investigation into generating levels with ChatGPT can help researchers to better understand the ability and limitations of LLMs.
-Talakat [96] is a bullet hell game engine which defines a domain-specific language for bullet hell levels, supporting the grammar evolution of bullet hell game content.
-Keiki [101] is another bullet hell type game engine which provides a native Python interface describing bullet hell barrage.
-While Talakat represents barrages as scripts, Keiki provides a function to convert barrage scripts into vector-based representations, making it convenient for ML-based content generators.
-IV-B2 3D Map & Level
+4.2.2 3D Map & Level
 3D maps and levels are closer to real-world scenarios but their generation is more complex.
@@
 Figure 6: Screenshots of cities generated by an entry [160] to the GDMC AI Settlement Generation Challenge in Minecraft [159] in 2023.
-IV-B3 Narrative
+4.2.3 Narrative
 AI techniques can be used to generate interactive fiction worlds with semantic coherence and fun.
@@
 The aim of Textworld is to provide an accessible API for RL agents [161] to understand natural language and then make decisions.
-IV-B4 Rhythm Chart
+4.2.4 Rhythm Chart
 Rhythm game is a special 2D game that aims to hit charts at the right time.
@@
 StepMania [142] is a free dance and rhythm game engine based on game Dance Dance Revolution that supports customised rhythm charts.
-StepMania has been researched in level generation i.e.
-chart generation [126], which is related to the music information retrieval field using long short-term memory (LSTM) and C-LSTM model.
+StepMania has been researched in level generation, i.e., chart generation [162], which is related to the music information retrieval field using long short-term memory (LSTM) and C-LSTM model.
 StepMania can be used in difficulty estimation [163] as well.
@@
 [143].
-IV-B5 Board & Card
+4.2.5 Board & Card
 Content generation for board and card games encompasses the rules, mechanism or deck building.
@@
 Magic: The Gathering is another commercial CCG, which has an open-source simulator called XMage for AI research including text design [167] and game playing [168].
-IV-C Other Resources for AI-involved Game Design
+4.3 Other Resources for AI-involved Game Design
 PhyDSL provides an accessible environment to design physics-based 2D games via code generation [146].
@@
 Although the codes of some platforms are not released, interactive interfaces are accessible to game designers, e.g., Baba is y’all [149] and Lode Encoder [150].
-IV-D Guidance for Designing Games
-We summarise the games and platforms for game design in Tab.
-III.
-A majority of game design platforms focus on content generation.
+4.4 Guidance for Designing Games
+Tab.
+3 summarises the games and platforms for game design, in which how AI can be utilised can be checked in column Type.
 Programming languages vary across the games and platforms, with half of them supporting Java.
@@
 For example, researchers can access Java programs in Python via Jpype [^3] https://github.com/jpype-project/jpype/.
+Some specific games, such as FightICE and VGC, can be used for studying game balance and skill discovery via parameter tuning techniques [170, 171, 172].
+A majority of games for design focus on procedurally generating content such as 2D and 3D levels.
+AI techniques including rule-based, search-based, ML-based and DL-based methods have been widely used for PCG using games listed in Tab.
+3 [22, 23, 5, 24, 25, 6, 26, 27].
+For example, bullet hell and maze level can be realised by grammar evolution [134, 173] and quality diversity search [174].
+Real-time generation of 2D levels of Super Mario Bros.
+represented by latent vectors also benefits from RL [175, 153].
+These games also serve as benchmarks for generative models like generative adversarial networks [176], diffusion models [177] and LLMs [32].
+In addition, supervised learning approaches have been applied to rhythm chart generation [162, 165] using StepMania and OSU!, as well as card generation [144] based on XMage.
+Notably, games for content generation like Procedural Zelda are also involved with adjusting game difficulty [178], and tuning parameters of games or agents [152, 179, 180].
 Beginners are recommended to try 2D level generation first since it is the most studied with massive platforms.
@@
 Researchers can directly choose the corresponding platform according to the interested content type with the help of Tab.
-III.
+3.
 FightingICE, AIGD2 and VGC are platforms that can be used for studying game tuning and balancing, but with different focus on fighting games, RTS games and Pokémon, respectively.
-V Discussion and Outlook
+5 Discussion and Outlook
 New platforms and games have been published over the last few years as ideal test-beds for AI research, including multi-agent RL, artificial general intelligence and creative AI.
 This section discusses the core challenges from games, our key findings and outlook.
-V-A Challenges from Games
+5.1 Challenges from Games
 Although originally designed for human entertainment, games have proven to be ideal test-beds for validating the capabilities of AI techniques, providing a range of challenges for AI to overcome.
-V-A1 Curse of High Dimensions
+5.1.1 Curse of High Dimensions
 Dimensions of games are involved with the number of game states, representation complexity of state and action spaces.
@@
 Furthermore, video games often provide players with a large and complex set of actions, leading to a combinatorial explosion of possible actions [182].
-V-A2 Poor Exploration
+5.1.2 Poor Exploration
 AI techniques such as RL usually need a large amount of interaction data to train policies.
@@
 Integrating the exploration signal into the reward function has been shown efficient to tackle the challenge, such as count-based exploration [122], intrinsic curiosity [184] and novelty search [185].
-While the above reward shaping may lead to the optimal policy variance, population-based exploration [140] and noisy exploration [136, 141] that perturb the policy parameter space present their advantage for the diverse exploration.
-V-A3 Partial Observability
+While the above reward shaping may lead to the optimal policy variance, population-based exploration [186] and noisy exploration [183, 187] that perturb the policy parameter space, present their advantage for the diverse exploration.
+5.1.3 Partial Observability
 While complete information is known in Atari [93] and Go [14] at a given state, partial observability is encountered in more complex video games such as ViZDoom [85].
@@
 Although agents can handle situations in front of them, unexpected occurrences such as monsters approaching from behind should also be considered.
-V-A4 Generalisation
-Although AI techniques such as deep RL have achieved a competitive level beyond humans in games including Go [5] and Atari [68], their generalisation ability is questionable since they are usually only trained on a given game or level.
+5.1.4 Generalisation
+Although AI techniques such as deep RL have achieved a competitive level beyond humans in games including Go [14] and Atari [93], their generalisation ability is questionable since they are usually only trained on a single game or level.
 The generalisation considers if the agent can still play well in some unseen games and levels without overfitting the training environment [189].
@@
 Procgen [102] also provides diverse game scenarios, which test the AI agents by randomly changing the initial position of the avatars.
-V-A5 Multi-agent Games
+5.1.5 Multi-agent Games
 It is common to see the participation of multiple agents or players in games such as tabletop games [55, 49], multiplayer online battle arena games [16], and RTS games [15].
@@
 In this combination, LLMs can understand the current game state and extract information, which enables an RL agent to make reasonable decisions.
-V-B AI for Games
+5.2 AI for Games
 The traditional developing process of games is long and complex, which burdens humans in designing and programming game content.
 Nowadays, AI can assist humans to improve games by testing and generating new content.
-V-B1 Game Testing by AI
+5.2.1 Game Testing by AI
 Game quality assurance is crucial in the gaming industry for assessing game content [193].
@@
 The critical challenge lies in testing updated content and maintaining game balance.
-Such studied [152, 153] have been conducted on multiple games such as VGC [35] and StarCraft II [6].
+Such studies [198, 170] have been conducted on multiple games such as VGC [54] and StarCraft II [15].
 Advanced testing methodologies are capable of adapting to the rapidly changing landscape of game development, which enhances the player experience without compromising game quality or balance.
-V-B2 Game Design by AI
+5.2.2 Game Design by AI
 In games, levels is usually the most fantastic part, which represents the effort and inspiration of game designers.
@@
 It is believed that there will be more AIs capable of generating complete games, marking the dawn of a new era in game development.
-V-C Games for AI
+5.3 Games for AI
 Beside the specific challenges such as high dimensions and partial observability, games can enhance AI by providing unique testing environments.
 Games not only expand AI capabilities, but also are regarded as preliminary experiments before applying AI to real-world scenarios.
-V-C1 Open-world Games for Artificial General Intelligence
+5.3.1 Open-world Games for Artificial General Intelligence
 General (video) game playing has been extensively studied using multi-game platforms [87, 95, 96, 127].
@@
 The diverse tasks such as farming, hunting and crafting in Minecraft enable AI to live as humans.
-This case is pushing the boundaries of AI towards achieving artificial general intelligence, as they mimic the complexity and unexpected of real-world tasks and decision-making.
-V-C2 Games as Reflection of Real-world Scenarios
+This case is pushing the boundaries of AI towards achieving artificial general intelligence, as it mimics the complexity and unexpected cases of real-world tasks and decision-making.
+5.3.2 Games as Reflection of Real-world Scenarios
 Games can simulate real-world scenarios and enable us to make new attempts of AI safely at a low cost and reflect challenges in real-world scenarios.
@@
 However, the codes of those commercial games are not released.
-V-C3 Emerging Environments for AI Creativity
+5.3.3 Emerging Environments for AI Creativity
 The environments collected in this paper indicate the research trend from intelligent decision-making to creative AI.
@@
 More open-source environments with diverse genres and content types, user-friendly APIs and documents, are expected to advance the research in AI creativity and human-AI co-creativity.
-V-D Under-explored Game Genres
-Although there are many available games, they share a large number of similar mechanisms with only differing in sprite styles.
+5.4 Under-explored Game Genres
+Although there are many available games, they share a large number of similar mechanisms with only differences in sprite styles.
 Many platforms for AI research are based on classic games like Atari [93].
-Although the simplified representation diminishes some engineering problems, this may not lead to a more delicate game mechanism thereby promoting AI research.
+Although the simplified representation diminishes some engineering problems, it is hard to construct more delicate game mechanisms.
 Only a few modern commercial games such as StarCraft II [15], and Dota 2 [16], provide accessible APIs or lightweight simulators.
@@
 [72], Angry Birds [130, 78] and Minecraft [82, 83].
-Though using classical games makes it convenient for research, it can prevent AI-involved game design from being used in practice.
-V-E Lower Barriers for Building New Games
-There are some professional game engines to build simulators for machine learning such as Unity [80], Unreal Engine [^5] https://github.com/zfw1226/gym-unrealcv and Godot [160].
+Though using classical games makes it convenient for research, it prevents AI-involved game design from being used in practice.
+5.5 Lower Barriers for Building New Games
+There are some professional game engines to build simulators for ML research such as Unity [106], Unreal Engine [^5] https://github.com/zfw1226/gym-unrealcv and Godot [205].
 Developing a domain-specific platform for a game is not easy.
-Some simulators can only be used in specific scenarios, and if the AI algorithm needs to be used in other situations, a new simulator may need to be built, which is time-consuming.
+Some simulators can only be used in specific scenarios, and if the AI algorithm needs to be used in other situations, building a new simulator may be necessary, which is time-consuming.
 One way to make simulators more versatile is by applying description languages.
@@
 With remote procedure call, algorithms do not need to be implemented with the same programming language as simulators do, which makes it possible to use general platforms for AI research, reducing engineering issues.
-V-F Large Language Models for Games
+5.6 Large Language Models for Games
 Recently, the emergence of LLMs such as ChatGPT [191] narrows the gap of game generation [157].
-Using natural languages, i.e., interacting with LLMs, everyone can create their own game.
+Using natural languages, i.e., interacting with LLMs, everyone can create their own games.
 Some other applications of LLMs for assisting game development such as level generation and dialogue generation also thrive.
@@
 The work of [30] validates the effectiveness of LLMs for generating Sokoban levels from the perspective of playability, novelty, diversity and accuracy.
-Instead of assisting game content generation, LLM itself can also be regarded as a part of games.
+Instead of assisting game content generation, LLMs themselves can also be regarded as a part of games.
 NPCs have been playing an essential role in games.
@@
 Kumaran et al.
-[18] propose SceneCraft based on LLM for narrative generation in a 3D game.
-In addition to dialogue, SceneCraft generates emotion and gestures of NPCs during the interactions with human players.
+[29] propose SceneCraft based on LLMs for narrative generation in a 3D game.
+In addition to dialogue, SceneCraft [29] generates emotion and gestures of NPCs during the interactions with human players.
 The work of [207] shows the possibility of developing software including games using multiple LLM agents.
@@
 [33] use GPT-4 to read a paper about game Spring, and then directly apply GPT-4 to play the game.
-According to [22], GPT-4 presents outstanding performance than advanced RL algorithms without training.
-It is widely believed that LLMs, with their remarkable capabilities and capacity for generalisation, are driving the future direction of game and game AI development.
-V-G Roles of Games in Interdisciplinary Research in AI
+According to [33], GPT-4 presents an outstanding performance than advanced RL algorithms without training.
+It is widely believed that LLMs, with their remarkable capabilities and capacity for generalisation, are driving the future direction of games and game AI development.
+5.7 Roles of Games in Interdisciplinary Research in AI
 Games not only play an important role in the development of AI, but also is an ideal tool for studying interdisciplinary research in AI and other computing-related areas, such as social intelligence of AI [208], affective computing [27], health and education [209, 210, 211], cognitive behavioural therapy for children [212].
@@
 Games have also been used as the mediums for educating AI [214].
-VI Conclusion
+6 Conclusion
 This paper reviews and categorises the publicly available games and game-based platforms for AI research.
-Growing research interests in AI for game testing, and AI for game design, games for artificial general intelligence, games for creative AI, have been induced.
-We present the games and game-based platforms for playing according to their characteristics, such as aims, the number of agents, observability, programming languages, and whether AI related competitions are held.
+Growing research interests in AI for game testing, and AI for game design, games for artificial general intelligence, multi-agent systems, and games for creative AI, have been induced.
+We present the games and game-based platforms for playing according to their characteristics, such as aims, the number of agents, observability, programming languages, and whether AI related competitions were held.
 Games and platforms for design are grouped by focused content type, such as 2D levels, 3D levels, narratives, rhythm charts, and cards.
@@
 (ii) What the games/platforms reviewed by this paper can contribute to the development of LLMs?
+Acknowledgement
+Authors would like to thank all anonymous reviewers for their careful review and insightful comments.